### v1

# Pointwise convergence of ergodic averages with Möbius weight

Abstract.
Let $(X,\nu,T)$ be a measure-preserving system, and let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients. We prove that, for any $f_{1},\ldots,f_{k}\in L^{\infty}(X)$, the Möbius-weighted polynomial multiple ergodic averages
$\frac{1}{N}\sum_{n\leq N}\mu(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)$
converge to 00 pointwise almost everywhere, provided that either: (i) $k\leq 3$ and $P_{i}$ are linear; (ii) $k$ is arbitrary and $P_{i}$ have distinct degrees; (iii) $k$ and $P_{i}$ are arbitrary, but $N$ is restricted to a certain explicit sequence of natural numbers. Specialising to $P_{1}(y)=y,P_{2}(y)=2y$, our results solve a problem of Frantzikinakis. For the proofs we establish some quantitative generalised von Neumann theorems for polynomial configurations that are of independent interest.
1. Introduction
Let $(X,\nu)$ be a probability space and $T\colon X\to X$ an invertible measure-preserving map, meaning that $\nu(T^{-1}(E))=\nu(E)$ for all measurable sets $E\subset X$. The triple $(X,\nu,T)$ is called a measure-preserving system. Given functions $f_{1},\ldots,f_{k}\in L^{\infty}(X)$ and polynomials $P_{1},\ldots,P_{k}\in\mathbb{Z}[y]$, we form the polynomial multiple ergodic averages
(1.1) $\frac{1}{N}\sum_{n\leq N}f_{1}(T^{P_{1}(n)})\cdots f_{k}(T^{P_{k}(n)}x),\quad x\in X.$
The convergence properties of these averages as $N\to\infty$ have been studied intensively. The question of $L^{2}(X)$ norm convergence of these averages was settled, after a series of partial results, by Host–Kra [16] and Leibman [20]. The question of pointwise convergence, however, remains open and is the subject of the celebrated Furstenberg–Bergelson–Leibman conjecture [2, Section 5.5]. Pointwise convergence has so far been established in two notable cases, namely the case where $k=1$ and the case where $k=2$ and $P_{1}$ is linear. The former follows from celebrated work of Bourgain [3], and the latter follows from another well known work of Bourgain [4] if $P_{2}$ is linear and from a recent breakthrough of Krause, Mirek and Tao [18] if $P_{2}$ has degree at least $2$.
Let $\mu$ be the Möbius function, defined as $\mu(n)=(-1)^{k}$ if $n$ is the product of $k$ distinct primes, and $\mu(n)=0$ if $n$ is divisible by the square of a prime. In this paper, we shall consider the Möbius-weighted polynomial multiple ergodic averages
$\frac{1}{N}\sum_{n\leq N}\mu(n)f_{1}(T^{P_{1}(n)})\cdots f_{k}(T^{P_{k}(n)}x),\quad x\in X.$
Based on the Möbius randomness principle (see [17, Section 13]), it is natural to conjecture the following.
Conjecture 1.1.
Let $k\in\mathbb{N}$, and let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients. Let $(X,\nu,T)$ be a measure-preserving system. Then, for any $f_{1},\ldots,f_{k}\in L^{\infty}(X)$, we have
(1.2) $\lim_{N\to\infty}\frac{1}{N}\sum_{n\leq N}\mu(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)=0$
for almost all $x\in X$.
The corresponding $L^{2}(X)$ norm convergence result was proven by Frantzikinakis and Host [8]. The only case we are aware of where Conjecture 1.1 was previously known is the case $k=1$ and $P_{1}$ linear, proven in [6, Proposition 3.1] (see also [8, Theorem C] for a generalisation of that result to other multiplicative functions).
Our first main theorem settles Conjecture 1.1 for $k=2$ in full, for $k=3$ if $P_{1},P_{2},P_{3}$ are linear, and for $k\geq 3$ if $P_{1},\ldots,P_{k}$ have distinct degrees. In fact, we get a polylogarithmic rate of converge to 00. These results go beyond what is currently known about the unweighted averages (1.1).
Theorem 1.2.
Let $k\in\mathbb{N}$, and let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients. Suppose that one of the following holds:
  1. (i)
    $k\leq 3$ and $P_{1},\ldots,P_{k}$ are linear polynomials;
  2. (ii)
    $P_{1},\ldots,P_{k}$ have pairwise distinct degrees.
Let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$. Let $(X,\nu,T)$ be a measure-preserving system. Then, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$ and $A\geq 0$, we have
$\lim_{N\to\infty}\frac{(\log N)^{A}}{N}\sum_{n\leq N}\mu(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)=0$
for almost all $x\in X$.
Let us make a few remarks about Theorem 1.2.
  1. (1)
    Specialising to $k=2$ and $P_{1}(y)=y,P_{2}(y)=2y$ (and $A=0$, $q_{1}=q_{2}=\infty$), Theorem 1.2 settles the Möbius case of Problem 12 of Frantzikinakis’ open problems survey [7] [^1] Frantzikinakis also asked about the convergence of the same multiple ergodic averages weighted by any multiplicative function $f\colon\mathbb{N}\to\mathbb{C}$, taking values in the unit disc, which has convergent means in the sense that $\lim_{N\to\infty}\mathbb{E}_{n\leq N}f(an+b)$ exists for all $a,b\in\mathbb{N}$. This general question remains open; see Remark 6.2.. This problem was also stated by Frantzikinakis and Host in [8].
  2. (2)
    In case (i), we can allow iterates of three different commuting transformations in the result; see Theorem 1.5 below.
  3. (3)
    It is likely that, at least for $k=2$, the region of $(q_{1},\ldots,q_{k})$ in the theorem could be improved to $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}\leq 1+\delta_{d}$ for some $\delta_{d}>0$, hence “breaking duality” in this problem. This is thanks to the $L^{p}$-improving estimates of Lacey [19] (see also [5]) and Han–Kovač–Lacey–Madrid–Yang [15]. See also [18, Section 11]. We leave the details to the interested reader.
  4. (4)
    The theorem continues to hold if we replace the Möbius function $\mu$ with the Liouville function $\lambda$ (defined by $\lambda(n)=(-1)^{\Omega(n)}$, where $\Omega(n)$ is the number of prime factors of $n$ with multiplicities); see the more general Theorem 6.1 along with Remark 5.3. Similarly, all the results in this paper regarding the Möbius function also hold for the Liouville function.
Our second main theorem shows that Conjecture 1.1 holds in general along a universal subsequence of $N$ (in fact, any sequence that grows fast enough).
Theorem 1.3.
Let $(N_{j})_{j\in\mathbb{N}}$ be any sequence satisfying $N_{j}\gg\exp(\exp(\exp((\log j)^{2})))$.
Let $k\in\mathbb{N}$, and let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients. Let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$. Let $(X,\nu,T)$ be a measure-preserving system. Then, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$, we have
$\lim_{j\to\infty}\frac{\exp((\log j)^{3/2})}{N_{j}}\sum_{n\leq N_{j}}\mu(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)=0$
for almost all $x\in X$.
We also show that Conjecture 1.1 holds if we replace the Cesàro average in (1.2) by a weaker double-logarithmic average. In the rest of the paper, for a function $F\colon[N]\to\mathbb{C}$ we define its double-logarithmic average up to $N$ as
(1.3) $\mathbb{E}_{n\leq N}^{\log\log}F(n)\coloneqq\left(\sum_{2\leq n\leq N}\frac{F(n)}{n\log n}\right)/\left(\sum_{2\leq n\leq N}\frac{1}{n\log n}\right).$
From partial summation one sees that, for any bounded function $F$, the existence of the limit $\lim_{N\to\infty}\frac{1}{N}\sum_{n\leq N}F(n)$ implies the existence of the limit $\lim_{N\to\infty}\mathbb{E}^{\log\log}_{n\leq N}F(n)$, but the converse is not true in general.
Theorem 1.4.
Let $k\in\mathbb{N}$, and let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients. Let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$. Let $(X,\nu,T)$ be a measure-preserving system. Then, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$, we have
$\lim_{N\to\infty}(\log\log N)^{c}\,\mathbb{E}_{n\leq N}^{\log\log}\mu(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)=0$
for almost all $x\in X$, for any $c>0$ small enough in terms of $P_{1},\ldots,P_{k},q_{1},\ldots,q_{k}$.
1.1. Ergodic averages with commuting transformations
Let a probability space $(X,\nu)$ and invertible, commuting, measure-preserving maps $T_{1},\ldots,T_{k}\colon X\to X$ be given. For any polynomials $P_{1},\ldots,P_{k}$ with integer coefficients and $f_{1},\ldots,f_{k}\in L^{\infty}(X)$, one can consider the more general polynomial ergodic averages with commuting transformations
(1.4) $\frac{1}{N}\sum_{n\leq N}f_{1}(T_{1}^{P_{1}(n)}x)\cdots f_{k}(T_{k}^{P_{k}(n)}x),\quad x\in X$
and their Möbius-weighted versions
(1.5) $\frac{1}{N}\sum_{n\leq N}\mu(n)f_{1}(T_{1}^{P_{1}(n)}x)\cdots f_{k}(T_{k}^{P_{k}(n)}x),\quad x\in X.$
The commuting case seems to be much more difficult, since in the unweighted case (1.4) pointwise convergence is not currently known even for $k=2$ and $P_{1},P_{2}$ linear (see Problem 19 of [7]). However, we mention that Walsh [28] proved $L^{2}(X)$ norm convergence of the averages (1.4) in a groundbreaking work, and the Furstenberg–Bergelson–Leibman conjecture states that pointwise convergence should hold even with commuting transformations.
Our next theorem states that for the Möbius averages (1.5) we have pointwise convergence in the case $k=3$ and $P_{1},P_{2},P_{3}$ linear.
Theorem 1.5.
Let $(X,\nu)$ be a probability space and let $T_{1},T_{2},T_{3}\colon X\to X$ be invertible, commuting, measure-preserving maps. Let $1<q_{1},q_{2},q_{3}\leq\infty$ satisfy $\frac{1}{q_{1}}+\frac{1}{q_{2}}+\frac{1}{q_{3}}<1$, and let $f_{1}\in L^{q_{1}}(X),f_{2}\in L^{q_{2}}(X),f_{3}\in L^{q_{3}}(X)$. Then, for any $A\geq 0$, we have
$\lim_{N\to\infty}\frac{(\log N)^{A}}{N}\sum_{n\leq N}\mu(n)f_{1}(T_{1}^{n}x)f_{2}(T_{2}^{n}x)f_{3}(T^{n}x)=0$
for almost all $x\in X$.
Naturally, this theorem implies Theorem 1.2(i) by restricting to $T_{1},T_{2},T_{3}$ being powers of the same transformation.
It seems likely that also Theorems 1.3 and 1.4 could be extended to the case of commuting transformations by extending Theorem 4.2 that goes into their proof to the case of multivariate functions (which could likely be done with a more complicated PET induction scheme). We leave the details of this extension to the interested reader.
1.2. Prime ergodic averages
The arguments presented in this paper are not limited to polynomial ergodic averages weighted by Möbius function, but instead apply to similar ergodic averages with any weight that satisfies certain Gowers uniformity assumptions as well as some upper bound assumptions that are easy to verify; see Theorems 6.16.46.6. In particular, thanks to the quantitative Gowers uniformity estimates for the von Mangoldt function in [22][26], these general theorems can be applied to reduce the problem of convergence of polynomial ergodic averages weighted by the primes to the problem of convergence of the same averages weighted by integers with no small prime factors, which is an easier problem (though still highly nontrivial). Pointwise convergence of polynomial multiple ergodic averages weighted by the primes will be studied in a future joint work with Krause, Mousavi and Tao.
1.3. Further applications of the proof method
A key ingredient in the proofs of our main theorems are some new polynomial generalised von Neumann theorems with quantitative dependencies that we establish in Section 4. These results are likely to have applications also to other problems than ergodic averages, in particular to quantitative asymptotics for the Möbius and von Mangoldt functions along polynomial progressions and to bounds for sets of integers lacking progressions of the form $x,x+P_{1}(p-1),\ldots x+P_{k}(p-1)$ with $p$ prime and with $P_{1},\ldots,P_{k}$ polynomials of distinct degrees with $P_{i}(0)=0$. Such applications will be investigated in future works.
1.4. Acknowledgements
The author thanks Nikos Frantzikinakis, Ben Krause, Sarah Peluse, Sean Prendiville and Terence Tao for helpful discussions and suggestions. The author was supported by funding from European Union’s Horizon Europe research and innovation programme under Marie Skłodowska-Curie grant agreement No 101058904.
2. Proof ideas
We now give an overview of the arguments used to prove Theorems 1.21.31.4, presenting the steps of the proof in a somewhat different order than in the actual proof and focusing on the case of functions $f_{i}\in L^{\infty}(X)$ for simplicity.
Let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients and with highest degree $d$. The starting point of the proofs of our pointwise convergence results is a lacunary subsequence trick combined with the Borel–Cantelli lemma and Markov’s inequality, which reduce matters to obtaining strong quantitative pointwise bounds for the averages
(2.1) $\left|\frac{1}{N^{d+1}}\sum_{m\leq N^{d}}\sum_{n\leq N}\mu(n)\phi(T^{m}x)f_{1}(T^{m+P_{1}(n)}x)\cdots f_{k}(T^{m+P_{k}(n)}x)\right|$
for $\nu$-almost all $x$ and all $\phi\in L^{\infty}(X)$. This reduction is presented in Section 6. For proving Theorem 1.2, we need bounds that save a large power of $(\log N)^{-1}$, whereas for proving Theorem 1.3 we need bounds that save a small power of $(\log\log N)^{-1}$ (and these suffice also for Theorem 1.4).
We then establish (in Section 4) polynomial generalised von Neumann theorems for counting operators of the form (2.1), which allow bounding the averages (2.1) in terms of the $U^{s}[N]$ Gowers norm of $\mu$ for some $s\in\mathbb{N}$, or some weaker norm such as the $u^{s}[N]$ norm (defined in Section 3). See in particular Theorems 4.1 and 4.2. Crucially, the bounds we obtain for (2.1) in terms of these norms of $\mu$ are quantitative with polynomial dependencies. These polynomial generalised von Neumann theorems are proven, depending on the case, either using the Peluse–Prendiville degree-lowering theory [23][24] or repeated use of van der Corput’s inequality coupled with the PET induction scheme.
After establishing the desired generalised von Neumann theorems, we can conclude the argument by applying the strongest known quantitative bounds for the $U^{s}[N]$ and $u^{s}[N]$ norms of the Möbius function (presented in Lemma 5.1). In the case of Theorem 1.2(ii), we use the fact that for any $s\in\mathbb{N}$ the $u^{s}[N]$ norm of $\mu$ has an arbitrary power of logarithm of cancellation (which follows from the Green–Tao Möbius and nilsequences result [13] but could also be proven by more classical means). In the case of Theorem 1.2(i), the proof is finished by using a recent quantitative $U^{4}[N]$ Gowers norm estimate for $\mu$ due to Leng [22] (which saves an arbitrary power of logarithm). For Theorems 1.3 and 1.4, in turn, we use a quantitative $U^{s}[N]$ Gowers norm estimate for $\mu$, due to Tao and the author [26] (which saves a small power of double-logarithm).
3. Notation and preliminaries
3.1. Asymptotic notation, indicators and averages
We use the Vinogradov and Landau asymptotic notations $\ll,\gg,o(\cdot),O(\cdot)$. Thus, we write $X\ll Y$, $X=O(Y)$ or $Y\gg X$ if there is a constant $C$ such that $|X|\leq CY$. We use $X\asymp Y$ to denote $X\ll Y\ll X$. We write $X=o(Y)$ as $N\to\infty$ if $|X|\leq c(N)Y$ for some function $c(N)\to 0$ as $N\to\infty$. If we add subscripts to these notations, then the implied constants can depend on these subscripts. Thus, for example $X\ll_{A}Y$ means that $|X|\leq C_{A}Y$ for some $C_{A}>0$ depending on $A$.
For a set $E$, we define the indicator function $1_{E}(x)$ as the function that equals to $1$ if $x\in E$ and equals to 00 otherwise. Similarly, if $P$ is a proposition, the expression $1_{P}$ equals to $1$ if $P$ is true and 00 if $P$ is false.
For a nonempty finite set $A$ and a function $f\colon A\to\mathbb{C}$, we define the averages
$\mathbb{E}_{a\in A}f(a)\coloneqq\frac{\sum_{a\in A}f(a)}{\sum_{a\in A}1}.$
We also use the double-logarithmic averages $\mathbb{E}^{\log\log}_{n\leq N}$, defined in (1.3).
For a real number $N\geq 1$, we denote $[N]\coloneqq\{n\in\mathbb{N}\colon\,\,n\leq N\}$. For integers $m,n$, we denote their greatest common divisor by $(m,n)$ and write $m\mid n^{\infty}$ to mean that $m\mid n^{k}$ for some natural number $k$. Unless otherwise specified, all our sums and averages run over the integers, with the exception that the variable $p$ is reserved for primes.
3.2. Gowers norms
For $s\in\mathbb{N}$ and a function $f\colon\mathbb{Z}\to\mathbb{C}$ with finite support, we define its unnormalised $U^{s}$ Gowers norm as
$\|f\|_{\widetilde{U}^{s}(\mathbb{Z})}\coloneqq\left(\sum_{x,h_{1},\ldots,h_{s}\in\mathbb{Z}}\prod_{\omega\in\{0,1\}^{s}}\mathcal{C}^{|\omega|}f(x+\omega\cdot(h_{1},\ldots,h_{s}))\right)^{1/2^{s}},$
where $\mathcal{C}(z)=\overline{z}$ is the complex conjugation operator and for a vector $(\omega_{1},\ldots,\omega_{s})$ we write $|\omega|\coloneqq|\omega_{1}|+\cdots+|\omega_{s}|$. For $N\geq 1$, we then define the $U^{s}[N]$ Gowers norm of a function $f\colon\mathbb{Z}\to\mathbb{C}$ as
$\|f\|_{U^{s}[N]}\coloneqq\frac{\|f1_{[N]}\|_{\widetilde{U}^{s}(\mathbb{Z})}}{\|1_{[N]}\|_{\widetilde{U}^{s}(\mathbb{Z})}}.$
As is well known (see for example [12, Appendix B]), for $s\geq 2$ the $U^{s}[N]$ norm is indeed a norm and for $s=1$ it is a seminorm, and the function $s\mapsto\|f\|_{U^{s}[N]}$ is increasing.
We observe the classical $U^{2}[N]$ inverse theorem: if $f\colon[N]\to\mathbb{C}$ satisfies $|f|\leq 1$ and $\delta\in(0,1)$, then
(3.1) $\|f\|_{U^{2}[N]}\geq\delta\implies\sup_{\alpha\in\mathbb{R}}|\mathbb{E}_{n\in[N]}f(n)e(-\alpha n)|\gg\delta^{2},$
where $e(x)\coloneqq e^{2\pi ix}$. This follows from the identity $\|f\|_{\widetilde{U}^{2}(\mathbb{Z})}=\int_{0}^{1}|\sum_{n\in\mathbb{Z}}f(n)e(-\alpha n)|^{4}\textnormal{ d}\alpha$ (which can be verified by expanding out the right-hand side) combined with Parseval’s identity.
For the $\widetilde{U}^{s}(\mathbb{Z})$ norms we have the Gowers–Cauchy–Schwarz inequality (see for example [26, (4.2)]), which states that, for any functions $(f_{\omega})_{\omega\in\{0,1\}^{s}}$ from $\mathbb{Z}$ to $\mathbb{C}$ with finite support, we have
(3.2) $\left|\sum_{x,h_{1},\ldots,h_{s}\in\mathbb{Z}}\prod_{\omega\in\{0,1\}^{s}}f_{\omega}(x+\omega\cdot(h_{1},\ldots,h_{s}))\right|\leq\prod_{\omega\in\{0,1\}^{s}}\|f_{\omega}\|_{\widetilde{U}^{s}(\mathbb{Z})}.$
We also define the $u^{s}[N]$ norm of a function $f\colon\mathbb{Z}\to\mathbb{C}$ as
(3.5) $\|f\|_{u^{s}[N]}\coloneqq\sup_{\begin{subarray}{c}P(y)\in\mathbb{R}[y]\\ \deg(P)\leq s-1\end{subarray}}\left|\mathbb{E}_{n\in[N]}f(n)e(-P(n))\right|.$
We remark that, as is well known, for $s=2$ the $u^{s}[N]$ and $U^{s}[N]$ norms are equivalent (up to polynomial losses), but for all $s\geq 3$ the $u^{s}[N]$ norm is weaker in the sense that the $U^{s}[N]$ norm controls the $u^{s}[N]$ norm but not vice versa; see [14][11, Section 4].
3.3. Van der Corput’s inequality
For $H\geq 1$, define the weight
$\mu_{H}(h)\coloneqq\frac{|\{(h_{1},h_{2})\in[H]^{2}\colon h_{1}-h_{2}=h\}|}{\lfloor H\rfloor^{2}},$
where $\lfloor x\rfloor$ is the floor of $x$ (This should not be confused with the Möbius function $\mu$). For an integer $h$, we define the differencing operator $\Delta_{h}$ by setting $\Delta_{h}f(x)=\overline{f(x+h)}f(x)$ for any $f\colon\mathbb{Z}\to\mathbb{C}$ and $x\in\mathbb{Z}$.
We will frequently use van der Corput’s inequality in the following form.
Lemma 3.1.
For any $N\geq H\geq 1$ and any function $f\colon\mathbb{Z}\to\mathbb{C}$ supported on $[N]$, we have
(3.6) $\left|\mathbb{E}_{n\in[N]}f(n)\right|^{2}\leq\frac{N+H}{\lfloor N\rfloor}\sum_{h\in\mathbb{Z}}\mu_{H}(h)\mathbb{E}_{n\in[N]}\Delta_{h}f(n).$
Proof.
See for example [25, Lemma 3.1]. ∎
3.4. Vinogradov’s Fourier expansion
For a real number $x$, we write $\|x\|$ for the distance from $x$ to the nearest integer(s).
We shall need a Fourier approximation for the indicator function of an interval that goes back to Vinogradov.
Lemma 3.2.
For any real numbers $-1/2\leq\alpha<\beta\leq 1/2$ and $\eta\in(0,\min\{1/2-\|\alpha\|,1/2-\|\beta\|,\|\alpha-\beta\|/2\})$, there exists a $1$-periodic function $g\colon\mathbb{R}\to[0,1]$ with the following properties.
  1. (1)
    $g(x)=1$ for $x\in[\alpha+\eta,\beta-\eta]$, $g(x)=0$ for $x\in[-1/2,1/2]\setminus[\alpha-\eta,\beta+\eta]$, and $0\leq g(x)\leq 1$ for all $x\in[-1/2,1/2]$.
  2. (2)
    For some $|c_{j}|\leq 10\eta$, we have the pointwise convergent Fourier representation
    $g(x)=\beta-\alpha-\eta+\sum_{|j|>0}c_{j}e(jx).$
  3. (3)
    For any $K\geq 1$, we have
    $\sum_{|j|>K}|c_{j}|\leq\frac{10\eta^{-1}}{K}.$
Proof.
This follows from [27, Lemma 12] (taking $r=1$ there). ∎
4. Polynomial generalised von Neumann theorems
In this section, we prove generalised von Neumann theorems for the weighted polynomial counting operators
(4.1) $\Lambda^{M,N}_{P_{1},\ldots,P_{k}}(\theta;f_{0},f_{1},\ldots,f_{k})\coloneqq\frac{1}{MN}\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\theta(n)f_{0}(m)f_{1}(m+P_{1}(n))\cdots f_{k}(m+P_{k}(n)),$
where $P_{1},\ldots,P_{k}\in\mathbb{Z}[y]$ and $\theta\colon[N]\to\mathbb{C}$ is a weight function (which we take in applications to be the Möbius function) and $f_{0},f_{1},\ldots,f_{k}\colon\mathbb{Z}\to\mathbb{C}$ are functions supported on $[-M,M]$ (with $M\asymp N^{\max_{j}(\deg P_{j})}$). Thus, we bound the expression (4.1) in terms of some Gowers norm (or related norm) of $\theta$ or $f_{0}$. It is important for the proofs of our main theorems that the obtained results are quantitative, with polynomial dependencies. The two main results of this section (which are of independent interest) are Theorems 4.1 and 4.2.
The first main result of this section states that if $P_{1},\ldots,P_{k}$ have distinct degrees, then the polynomial counting operator (4.1) is bounded in terms of the $u^{s}[N]$ norm of the weight $\theta$ for some $s$, with polynomial dependencies. It is important to have the $u^{s}[N]$ norm rather than the $U^{s}[N]$ norm here, since for the former we have much better quantitative bounds in the case of the Möbius function.
Theorem 4.1 (A polynomial generalised von Neumann theorem with $u^{s}$ control).
Let $d,k\in\mathbb{N}$ and $C\geq 1$. Let $P_{1},\ldots,P_{k}$ be a polynomials with integer coefficients satisfying $\deg P_{1}<\deg P_{2}<\cdots<\deg P_{k}=d$. Let $N\geq 1$, and let $f_{0},\ldots,f_{k}\colon\mathbb{Z}\to\mathbb{C}$ be functions supported on $[-CN^{d},CN^{d}]$ with $|f_{i}|\leq 1$ for all $0\leq i\leq k$, and let $\theta\colon[N]\to\mathbb{C}$ be a function with $|\theta|\leq 1$. Then, for some $11\leq K\ll_{d}$ 1, we have
$\left|\frac{1}{N^{d+1}}\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\theta(n)f_{0}(m)f_{1}(m+P_{1}(n))\cdots f_{k}(m+P_{k}(n))\right|\ll_{C,P_{1},\ldots,P_{k}}(N^{-1}+\|\theta\|_{u^{d+1}[N]})^{1/K}.$
The proof of Theorem 4.1 is given in the next three subsections. In Subsection 4.1, we show that $f_{0},f_{1}$ can be assumed to be “locally linear phase functions” in a suitable sense. In Subsection 4.2, we prove the $k=1$ case of the theorem using the circle method; this works as a base case for the proof which is by induction on $k$. Finally, in Subsection 4.3, we use the conclusions of the preceding subsections together with an iterative argument for the function $f_{1}$ to conclude the proof.
The second main result of this section is that, for any polynomials $P_{1},\ldots,P_{k}$, the polynomial counting operator (4.1) is always bounded by some $U^{s}[N]$ norm of the weight $\theta$, with linear dependence on the Gowers norm. In what follows, for any finite nonempty collection $\mathcal{Q}$ of polynomials with integer coefficients, we define its degree $\deg\mathcal{Q}$ as the largest of the degrees of the polynomials in $\mathcal{Q}$.
Theorem 4.2 (A polynomial generalised von Neumann theorem with $U^{s}$ control).
Let $d\in\mathbb{N}$ and $C\geq 1$. Let $N\geq 1$, and let $\theta\colon[N]\to\mathbb{C}$ be a function. Let $\mathcal{Q}$ be a finite collection of polynomials with integer coefficients satisfying $\deg\mathcal{Q}=d$ and $Q([N])\subset[-CN^{d},CN^{d}]$ for all $Q\in\mathcal{Q}$. For each $Q\in\mathcal{Q}$ let $f_{Q}\colon\mathbb{Z}\to\mathbb{C}$ be a function supported on $[-CN^{d},CN^{d}]$ with $|f_{Q}|\leq 1$. Then, for some natural number $s\ll_{|\mathcal{Q}|,\deg\mathcal{Q}}1$, we have
(4.2) $\left|\frac{1}{N^{d+1}}\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\theta(n)\prod_{Q\in\mathcal{Q}}f_{Q}(m+Q(n))\right|\ll_{|\mathcal{Q}|,\deg\mathcal{Q},C}\|\theta\|_{U^{s}[N]}.$
The proof of Theorem 4.2 is based on the PET induction scheme and is given in Subsection 4.4. We also present there a multidimensional version of the special case where $\deg\mathcal{Q}=1$ (Lemma 4.7); this will be needed for the proof of Theorem 1.5.
4.1. Transferring to locally linear functions
[…]
Proposition 4.3 (Reduction to locally linear phases).
Let $C\geq 1$, $d,k\in\mathbb{N}$, and let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients and with $\deg P_{1}<\cdots<\deg P_{k}=d$. Let $N\geq 1$, and let $f_{0},f_{1},\ldots,f_{k}\colon\mathbb{Z}\to\mathbb{C}$ be functions supported on $[-CN^{d},CN^{d}]$ and with $|f_{i}|\leq 1$ for all $0\leq i\leq k$. Let $\theta\colon[N]\to\mathbb{C}$ be a function with $|\theta|\leq 1$. Then, for some $11\leq K\ll_{d}$ 1, we have
$\left|\frac{1}{N^{d+1}}\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\theta(n)f_{0}(m)f_{1}(m+P_{1}(n))\cdots f_{k}(m+P_{k}(n))\right|$
$\quad\ll_{C,P_{1},\ldots,P_{k}}\left(N^{-1}+\left|\frac{1}{N^{d+1}}\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\theta(n)\phi_{0}(m)\phi_{1}(m+P_{1}(n))\cdots f_{k}(m+P_{k}(n))\right|\right)^{1/K}$
for some locally linear phase functions $\phi_{0},\phi_{1}$ of resolution $\gg_{C,P_{1},\ldots,P_{k}}\delta^{O_{d}(1)}N^{\deg P_{1}}$. Moreover, we may assume that the spectra of $\phi_{0},\phi_{1}$ belong to $\frac{1}{N^{\deg P_{1}}}\mathbb{Z}$.
Proposition 4.3 may be compared with, and is motivated by, the work of Peluse and Prendiville [24, Theorem 1.5], where in the case $k=2$, $\theta\equiv 1$ and $P_{1}(y)=y,P_{2}(y)=y^{2}$, it is proved that $f_{0},f_{1},f_{2}$ can be replaced more strongly with major arc locally linear phase functions. In the more general setup of Proposition 4.3, it is not possible to reduce to major arc locally linear phase functions.
For the proof of Proposition 4.3, we need Peluse’s inverse theorem.
Lemma 4.4 (Peluse’s inverse theorem).
Let $k,d_{1},d\in\mathbb{N}$ and $C\geq 1$. Let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients satisfying $P_{i}(0)=0$ for all $1\leq i\leq k$ and $d_{1}=\deg P_{1}<\ldots<\deg P_{k}=d$, and with all the coefficients of the polynomials $P_{i}$ being bounded by $C$ in modulus.
Let $N\geq 1$ and $\delta\in(0,1/2)$. Let $f_{1},\ldots,f_{k}\colon\mathbb{Z}\to\mathbb{C}$ be functions supported on $[-CN^{d},CN^{d}]$ with $|f_{i}|\leq 1$ for all $0\leq i\leq k$. Then there exists $11\leq B\ll_{d}$ 1 such that for either $j\in\{0,1\}$ we have
(4.3) $\begin{split}&\left|\frac{1}{N^{d+1}}\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}f_{0}(m)f_{1}(m+P_{1}(n))\cdots f_{k}(m+P_{k}(n))\right|\\ &\quad\ll_{C,d}\delta+\left(N^{-1}+\max_{\begin{subarray}{c}q\in[\delta^{-B}]\\ N^{\prime}\in[\delta^{B}N^{d_{1}},N^{d_{1}}]\end{subarray}}\frac{1}{N^{d}}\sum_{m\in\mathbb{Z}}\left|\frac{1}{N^{\prime}}\sum_{n\in[N^{\prime}]}f_{j}(m+qn)\right|\right)^{1/B}.\end{split}$
Proof.
This will follow from [23, Theorem 3.3] after some reductions. It suffices to show that for each $j\in\{0,1\}$ there exists $11\leq B\ll_{d}$ 1 such that (4.3) holds, since we may increase $B$ if necessary.
Suppose first that $j=0$. Using the notation (4.1), let
(4.4) $\eta\coloneqq|\Lambda^{CN^{d},N}_{P_{1},\ldots,P_{k}}(1;f_{0},\ldots,f_{k})|.$
Then $C\eta$ is equal to the left-hand side of (4.3). We may clearly assume that $\eta\geq\delta$ and that $\delta\leq 1/L$ for any given constant $L=L_{C,d}$. We may further assume that $N\geq\delta^{-K}\geq\eta^{-K}$ for any given constant $K=K_{d}$, since otherwise by taking $B=Kd_{1}$ the claim follows (with $q=1$ in (4.3)) from the crude triangle inequality bound
$|\Lambda^{CN^{d},N}_{P_{1},\ldots,P_{k}}(1;f_{0},\ldots,f_{k})|\leq\frac{1}{CN^{d}}\sum_{m\in[-CN^{d},CN^{d}]}|f_{0}(m)|.$
Now, applying [^2] In  [23, Theorem 3.3], the functions $f_{i}$ are assumed to be supported on $[CN^{d}]$ instead of $[-CN^{d},CN^{d}]$, but this makes no difference in the argument. [23, Theorem 3.3] (with $\eta$ in place of $\delta$) we see that there exists some $11\leq B\ll_{d}$ 1 such that
$\max_{\begin{subarray}{c}q\in[\delta^{-B}]\\ N^{\prime}\in[\delta^{B}N^{d_{1}},N^{d_{1}}]\end{subarray}}\frac{1}{N^{d}}\sum_{m\in\mathbb{Z}}\left|\frac{1}{N^{\prime}}\sum_{n\in[N^{\prime}]}f_{j}(m+qn)\right|\gg_{C,d}\eta^{B},$
which in view of (4.4) implies the claim.
Suppose then that $j=1$. Then, making the change of variables $m^{\prime}=m+P_{1}(n)$ in (4.1), we see that
$\Lambda^{CN^{d},N}_{P_{1},\ldots,P_{k}}(1;f_{0},f_{1},f_{2},\ldots,f_{k})=\Lambda^{CN^{d},N}_{-P_{1},P_{2}-P_{1},\ldots,P_{k}-P_{1}}(1;f_{1},f_{0},f_{2},\ldots,f_{k})$
Now the claim follows from the $j=0$ case handled above. ∎
Proof of Proposition 4.3.
We begin with a few reductions. Firstly, we may extend $\theta$ to a function on $\mathbb{Z}$ by setting it equal to 00 outside $[N]$. Secondly, we may assume that $P_{i}(0)=0$ for $1\leq i\leq k$ by translating the functions $f_{i}$ if necessary. Thirdly, we may assume that $C$ is large enough in terms of $P_{1},\ldots,P_{k}$ so that
$\max_{1\leq i\leq k}\max_{n\in[N]}|P_{i}(n)|\leq CN^{d}.$
Let
(4.5) $\eta\coloneqq|\Lambda_{P_{1},\ldots,P_{k}}^{CN^{d},N}(\theta;f_{0},f_{1},\ldots,f_{k})|.$
We may assume that $N\geq\eta^{-K}$ for any given constant $K$ depending on $d$, as otherwise the claim readily follows.
We first wish to replace $f_{0}$ with a locally linear phase function. For $m\in\mathbb{Z}$, define the first dual function
$F(m)\coloneqq\mathbb{E}_{n\in[N]}\theta(n)f_{1}(m+P_{1}(n))\cdots f_{k}(m+P_{k}(n)).$
Then by (4.5) we have
$\left|\sum_{m\in\mathbb{Z}}f_{0}(m)F(m)\right|\gg_{C}\eta N^{d},$
so by the Cauchy–Schwarz inequality we get
(4.6) $\left|\sum_{m\in\mathbb{Z}}\overline{F}(m)F(m)\right|\gg_{C}\eta^{2}N^{d}.$
By the definition of $F$, (4.6) expands out as
$\left|\sum_{m\in\mathbb{Z}}\mathbb{E}_{n\in[N]}\theta(n)\overline{F}(m)f_{1}(m+P_{1}(n))\cdots f_{k}(m+P_{k}(n))\right|\gg_{C}\eta^{2}N^{d}.$
Denote $d_{1}\coloneqq\deg P_{1}$. Applying Cauchy–Schwarz and van der Corput’s inequality (Lemma 3.1), this implies that
$\left|\sum_{h\in[-N^{d_{1}},N^{d_{1}}]}\mu_{N^{d_{1}}}(h)\mathbb{E}_{n\in[N]}\sum_{m\in\mathbb{Z}}\Delta_{h}\overline{F}(m)\Delta_{h}f_{1}(m+P_{1}(n))\cdots\Delta_{h}f_{k}(m+P_{k}(n))\right|\gg_{C}\eta^{4}N^{d}.$
Noting that $N^{d_{1}}|\mu_{N^{d_{1}}}(h)|\ll 1_{|h|\leq N^{d_{1}}}$, from the triangle inequality and the pigeonhole principle we now see that
(4.7) $\left|\mathbb{E}_{n\in[N]}\sum_{m\in\mathbb{Z}}\Delta_{h}\overline{F}(m)\Delta_{h}f_{1}(m+P_{1}(n))\cdots\Delta_{h}f_{k}(m+P_{k}(n))\right|\gg_{C}\eta^{4}N^{d}$
for $\gg_{C}\eta^{4}N^{d_{1}}$ integers $h\in[-N^{d_{1}},N^{d_{1}}]$.
Applying Lemma 4.4 and the pigeonhole principle, from (4.7) we conclude that there exists a constant $11\leq B\ll_{d}$ 1 and an integer $1\leq q\ll_{C,P_{1},\ldots,P_{k}}\eta^{-B}$ such that
(4.8) $\mathbb{E}_{h\in[-N^{d_{1}},N^{d_{1}}]}\max_{M\in[\eta^{B}N^{d_{1}},N^{d_{1}}]}\mathbb{E}_{m\in[-2CN^{d},2CN^{d}]}\left|\mathbb{E}_{y\in[qM]}1_{y\equiv 0\ (\mathrm{mod}\ q)}\Delta_{h}F(m+y)\right|\gg_{C,P_{1},\ldots,P_{k}}\eta^{B}.$
Using the simple bound
(4.9) $\mathbb{E}_{m\in[-X,X]}|\mathbb{E}_{y\in[Y]}a(m+y)|\ll\mathbb{E}_{m\in[-X,X]}|\mathbb{E}_{y\in[Y^{\prime}]}a(m+y)|+Y^{\prime}/Y+Y/X,$
valid for any bounded sequence $a\colon\mathbb{Z}\to\mathbb{C}$ and $1\leq Y^{\prime}\leq Y\leq X$, and setting
$N^{\prime}=\eta^{4B+2}N^{d_{1}},$
from (4.8) we deduce that
$\mathbb{E}_{h\in[-N^{d_{1}},N^{d_{1}}]}\mathbb{E}_{m\in[-2CN^{d},2CN^{d}]}\left|\mathbb{E}_{y\in[N^{\prime}]}1_{y\equiv 0\ (\mathrm{mod}\ q)}\Delta_{h}F(m+y)\right|\gg_{C,P_{1},\ldots,P_{k}}\eta^{B}.$
Splitting the $h$ average into intervals of length $2N^{\prime}$ and applying the pigeonhole principle, we see that there exists some integer $|\ell|\leq N^{d_{1}}$ such that
$\mathbb{E}_{h\in[-N^{\prime},N^{\prime}]}\mathbb{E}_{m\in[-2CN^{d},2CN^{d}]}\left|\mathbb{E}_{y\in[N^{\prime}]}1_{y\equiv 0\ (\mathrm{mod}\ q)}\Delta_{h+\ell}F(m+y)\right|\gg_{C,P_{1},\ldots,P_{k}}\eta^{B}.$
From Cauchy–Schwarz and van der Corput’s inequality, we then see that
(4.10) $\begin{split}&\mathbb{E}_{h\in[-N^{\prime},N^{\prime}]}\mathbb{E}_{h^{\prime}\in[-N^{\prime},N^{\prime}]}\lfloor N^{\prime}\rfloor\mu_{N^{\prime}}(h^{\prime})\mathbb{E}_{m\in[-2CN^{d},2CN^{d}]}\mathbb{E}_{y\in[N^{\prime}]}1_{y,y+h^{\prime}\equiv 0\ (\mathrm{mod}\ q)}\Delta_{h+\ell}\Delta_{h^{\prime}}F(m+y)\\ &\quad\gg_{C,P_{1},\ldots,P_{k}}\eta^{2B}.\end{split}$
We wish to remove the weight $\mu_{N^{\prime}}(h^{\prime})$ from (4.10). To this end, we note the easily verified Fourier expansion
[…]
for $x\in[-1,1]$, which allows us to write for $|h^{\prime}|\leq N$ the expansion
[…]
Substituting this to (4.10) and expanding $1_{y\equiv 0\ (\mathrm{mod}\ q)}=\sum_{0\leq r<q}e(ry/q)$ and using $e(\xi h^{\prime})=e(\xi(y+h+h^{\prime}))e(-\xi(y+h))$ and $e(\xi y)=e(\xi(y+h))e(\xi(y+h^{\prime}))e(-\xi(y+h+h^{\prime}))$, we see that
$\mathbb{E}_{h\in[-N^{\prime},N^{\prime}]}\mathbb{E}_{h^{\prime}\in[-N^{\prime},N^{\prime}]}\mathbb{E}_{m\in[-2CN^{d},2CN^{d}]}\mathbb{E}_{y\in[N^{\prime}]}\prod_{\omega\in\{0,1\}^{2}}F_{\omega}(m+y+\omega\cdot(h,h^{\prime}))$
$\quad\gg_{C,P_{1},\ldots,P_{k}}\eta^{2B},$
where $F_{(0,0)}=F$ and $|F_{\omega}(x)|\ll_{C}1$ for all $x,\omega$. Hence, by the Gowers–Cauchy–Schwarz inequality (3.2), we find
$\mathbb{E}_{m\in[-2CN^{d},2CN^{d}]}\|F\|_{U^{2}[m,m+N^{\prime}]}\gg_{C,P_{1},\ldots,P_{k}}\eta^{2B}.$
Applying the pigeonhole principle, we deduce that
(4.11) $\|F\|_{U^{2}[m,m+N^{\prime}]}\gg_{C,P_{1},\ldots,P_{k}}\eta^{2B}$
for $\gg\eta^{2B}N^{d}/N^{\prime}$ integers $m\in N^{\prime}\mathbb{Z}\cap[-CN^{d},CN^{d}]$.
Note that for any complex number $z$ we have $|z|\leq 10\max_{j\in\{0,1,2\}}\{\textnormal{Re}(e(j/3)z)\}$. For $j\in\{0,1,2\}$, let $\mathcal{M}_{j}$ be the set of $m\in N^{\prime}\mathbb{Z}\cap[-CN^{d},CN^{d}]$ for which
(4.12) $\sup_{\alpha\in[0,1]}\textnormal{Re}\left(e\left(\frac{j}{3}\right)\mathbb{E}_{x\in[m,m+N^{\prime}]}F(x)e(\alpha x)\right)\geq\eta^{4B+1}.$
Then, by the pigeonhole principle, the $U^{2}[N^{\prime}]$ inverse theorem (3.1) and (4.11), we have $|\mathcal{M}_{j_{0}}|\gg\eta^{2B}N^{d}/N^{\prime}$ for some $j_{0}\in\{0,1,2\}$. For any $m\in\mathcal{M}_{j_{0}}$, let $\alpha_{m}^{\prime}\in\frac{1}{N^{\deg P_{1}}}\mathbb{Z}$ be a point where the supremum in (4.12) is attained, and let $\alpha_{m}$ be an element of $\frac{1}{N^{d_{1}}}\mathbb{Z}$ nearest to $\alpha_{m}^{\prime}$. Recalling the definition of $N^{\prime}$, we have
$\textnormal{Re}\left(e\left(\frac{j_{0}}{3}\right)\mathbb{E}_{x\in[m,m+N^{\prime}]}F(x)e(\alpha_{m}x)\right)\gg_{C,P_{1},\ldots,P_{k}}\eta^{4B+1}.$
For $m\in N^{\prime}\mathbb{Z}\setminus\mathcal{M}_{j_{0}}$, note that by Parseval’s identity there is some $\alpha_{m}\in\frac{1}{N^{d_{1}}}\mathbb{Z}$ for which
$\left|\sum_{x\in[m,m+N^{\prime}]}F(x)e(\alpha_{m}x)\right|\ll_{C}(N^{\prime})^{1/2}.$
Now, extend the definition of $\alpha_{m}$ to all of $\mathbb{Z}$ by letting $\alpha_{m}=\alpha_{m^{\prime}}$, where $m^{\prime}$ is the largest element of $N^{\prime}\mathbb{Z}$ that is at most $m$. Then, define the locally linear phase function $\phi_{0}(m)=e(\alpha_{m}m)$, which has resolution $N^{\prime}$. We now have
$\left|\sum_{x\in[-CN^{d},CN^{d}]}F(x)\phi_{0}(x)\right|\gg_{C,P_{1},\ldots,P_{k}}\eta^{2B}\cdot\eta^{4B+1}$
Recalling the definition of $F$, we conclude that
(4.13) $\left|\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\theta(n)\phi_{0}(m)f_{1}(m+P_{1}(n))\cdots f_{k}(m+P_{k}(n))\right|\gg_{C,P_{1},\ldots,P_{k}}\eta^{B^{\prime}}N^{d+1},$
where $B^{\prime}=6B+1$.
We proceed to replace also $f_{1}$ with a locally linear phase function. For $m\in\mathbb{Z}$, define the second dual function as
$G(m)\coloneqq\mathbb{E}_{n\in[N]}\theta(n)\phi_{0}(m-P_{1}(n))f_{2}(m+P_{2}(n)-P_{1}(n))\cdots f_{k}(m+P_{k}(n)-P_{1}(n))$
Making a change of variables, from (4.13) it follows that
$\left|\sum_{m\in\mathbb{Z}}f_{1}(m)G(m)\right|\gg_{C,P_{1},\ldots,P_{k}}\eta^{B^{\prime}}N^{d}.$
Arguing verbatim as above, we deduce that there exists a locally linear phase function $\phi_{1}\colon\mathbb{Z}\to\mathbb{C}$ of resolution $N^{\prime\prime}=\eta^{B^{\prime\prime}}N^{d_{1}}$ and with spectrum in $\frac{1}{N^{d_{1}}}\mathbb{Z}$ such that
$\left|\sum_{m\in\mathbb{Z}}G(m)\phi_{1}(m)\right|\gg_{C,P_{1},\ldots,P_{k}}\eta^{B^{\prime\prime}}N^{d},$
where $B^{\prime\prime}=6B+1^{\prime}$.
Recalling the definition of $G$, this means that
$\left|\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\theta(n)\phi_{0}(m)\phi_{1}(m+P_{1}(n))f_{2}(m+P_{2}(n))\cdots f_{k}(m+P_{k}(n))\right|\gg_{C,P_{1},\ldots,P_{k}}\eta^{B^{\prime\prime}}N^{d+1}.$
This gives the desired claim. ∎
4.2. A circle method bound
The proof of Theorem 4.1 will proceed by induction on $k$, so we first need to bound the weighted averages (4.1) with $k=1$. These averages can be controlled simply by using classical Fourier analysis.
Lemma 4.5.
Let $d\in\mathbb{N}$ and $C\geq 1$. Let $P$ be a polynomial of degree $d$ with integer coefficients. Let $N\geq 1$, and let $f_{0},f_{1}\colon\mathbb{Z}\to\mathbb{C}$ be functions supported on $[-CN^{d},CN^{d}]$ with $|f_{i}|\leq 1$ for both $i\in\{0,1\}$, and let $\theta\colon[N]\to\mathbb{C}$ be a function. Then we have
(4.14) $\left|\frac{1}{N^{d+1}}\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\theta(n)f_{0}(m)f_{1}(m+P(n))\right|\ll_{C,P}\|\theta\|_{u^{d+1}[N]}.$
Proof.
Let $C^{\prime}\geq 1$ (depending on $C,P$) be such that $\max_{n\in[N]}|P(n)|\leq(C^{\prime}-C)N^{d}$. Then, by the orthogonality of characters, the left-hand side of (4.14) without absolute values equals to
$\frac{1}{N^{d+1}}\sum_{|a|\leq(C_{P}+C)N^{d}}\sum_{|m|\leq CN^{d}}\sum_{n\in[N]}\theta(n)f_{0}(m)f_{1}(a)1_{a=m+P(n)}$
$\quad=\frac{1}{N^{d+1}}\int_{0}^{1}\left(\sum_{|a|\leq C^{\prime}N^{d}}f_{1}(a)e(\xi a)\right)\left(\sum_{|m|\leq CN^{d}}f_{0}(m)e(-\xi m)\right)\left(\sum_{n\in[N]}\theta(n)e(-\xi P(n))\right)\textnormal{ d}\xi.$
Now the claim follows by bounding the exponential sum involving $\theta$ pointwise by $N\|\theta\|_{u^{d+1}[N]}$ and by using Cauhcy–Schwarz and Parseval’s identity to the remaining two exponential sums. ∎
4.3. Proof of Theorem 4.1
We are now ready to prove the claimed estimate for (4.1) for distinct degree polynomials.
Proof of Theorem 4.1.
We use induction on $k$. The base case $k=1$ follows from Lemma 4.5. Suppose that the case $k-1$ has been proven for some $k\geq 2$, and consider the case $k$.
Step 1: Reduction to locally linear phase functions. Let $\delta\coloneqq|\Lambda_{P_{1},\ldots,P_{k}}^{CN^{d},N}(\theta;f_{0},f_{1},\ldots,f_{k})|$. We may assume that $1/K^{\prime\prime}\geq\delta\geq N^{-1/K^{\prime}}$ for any large constants $K^{\prime}=K^{\prime}_{d}$ and $K^{\prime\prime}=K^{\prime\prime}_{C,P_{1},\ldots,P_{k}}$, as otherwise there is nothing to prove. By Proposition 4.3, there exist $C_{d}\geq 1$ and locally linear phase functions $\phi_{0},\phi_{1}\colon\mathbb{Z}\to\mathbb{C}$ of resolution $\geq N^{\prime}$ for some $N^{\prime}\gg_{C,P_{1},\ldots,P_{k}}\delta^{C_{d}}N^{\deg P_{1}}$, and with the spectra of $\phi_{0},\phi_{1}$ belonging to $\frac{1}{N^{\deg P_{1}}}\mathbb{Z}$, such that
(4.15) $|\Lambda_{P_{1},\ldots,P_{k}}^{CN^{d},N}(\theta;\phi_{0},\phi_{1},f_{2},\ldots,f_{k})|\geq\delta^{C_{d}}.$
Step 2: An iteration for the locally linear phase function. We can write $\phi_{1}(m)=e(\alpha_{m}m)$ for some $\alpha_{m}\in\frac{1}{N^{\deg P_{1}}}\mathbb{Z}$, with $m\mapsto\alpha_{m}$ is being constant on the the intervals $[jN^{\prime}+a,(j+1)N^{\prime}+a)$ for some $a\in[N^{\prime}]$ and all $j\in\mathbb{Z}$. For any set $S\subset\mathbb{R}$, write $\phi_{1,S}(m)\coloneqq\phi_{1}(m)1_{\alpha_{m}\not\in S}$.
Claim. If $C_{d}^{\prime}$ and $K^{\prime}_{d}$ are large, the following holds. For any $\eta\in(0,N^{-1/K_{d}^{\prime}})$ and any finite (possibly empty) set $S$, if
(4.16) $|\Lambda_{P_{1},\ldots,P_{k}}^{CN^{d},N}(\theta;\phi_{0},\phi_{1,S},f_{2},\ldots,f_{k})|\geq\eta,$
then either
(4.17) $|\Lambda_{P_{1},\ldots,P_{k}}^{CN^{d},N}(\theta;\phi_{0},\phi_{1}-\phi_{1,\{\alpha\}},f_{2},\ldots,f_{k})|\geq\eta^{2C_{d}^{\prime}}$
or there exists $\alpha\not\in S$ such that $\alpha_{m}=\alpha$ for $\geq\eta^{C_{d}^{\prime}}N^{d}$ integers $m\in[-CN^{d},CN^{d}]$ and
(4.18) $|\Lambda_{P_{1},\ldots,P_{k}}^{CN^{d},N}(\theta;\phi_{0},\phi_{1,S\cup\{\alpha\}},f_{2},\ldots,f_{k})|\geq\eta-\eta^{2C_{d}^{\prime}}.$
For proving this claim, we first apply van der Corput’s inequality (Lemma (3.1)) to (4.16) to conclude that there is a set $\mathcal{H}\subset[-CN^{d},CN^{d}]$ of size $\gg\eta^{2}N^{d}$ such that for $h\in\mathcal{H}$ we have
$|\Lambda_{P_{1},\ldots,P_{k}}^{CN^{d},N}(1;\Delta_{h}\phi_{0},\Delta_{h}\phi_{1,S},\Delta_{h}f_{2},\ldots,\Delta_{h}f_{k})|\gg\eta^{2}.$
From Lemma 4.4 and the pigeonhole principle, we now conclude that for some constant $B_{d}\geq 1$, some integer $1\leq q\leq\eta^{-B_{d}}$ and some $M\in[\eta^{B_{d}}N^{\deg P_{1}},N^{\deg P_{1}}]$, we have
$\mathbb{E}_{|x|\leq 2CN^{d}}|\mathbb{E}_{m\in[M]}\overline{\phi_{1,S}}(x+qm+h)\phi_{1,S}(x+qm)|\gg_{C,P_{1},\ldots,P_{k}}\eta^{B_{d}}$
for $\gg_{C,P_{1},\ldots,P_{k}}\eta^{B_{d}}N^{d}$ integers $h\in\mathcal{H}$. Let $\mathcal{H}^{\prime}$ be the set of such $h$. By (4.9) we also have
(4.19) $\mathbb{E}_{|x|\leq 2CN^{d}}|\mathbb{E}_{m\in[M^{\prime}]}\overline{\phi_{1,S}}(x+qm+h)\phi_{1,S}(x+qm)|\gg_{C,P_{1},\ldots,P_{k}}\eta^{B_{d}}$
for $h\in\mathcal{H}^{\prime}$, where $M^{\prime}=\eta^{4B_{d}+2}N^{\prime}$.
From the pigeonhole principle, we see that there exist $h_{0}\in[N^{\prime}]$ and $\mathcal{H}^{\prime\prime}\subset\mathcal{H}^{\prime}$ with |′′$|\mathcal{H}^{\prime\prime}|\gg_{C,P_{1},\ldots,P_{k}}\eta^{4B_{d}}N^{d}$ such that $\|(h-h_{0})/N^{\prime}\|\leq\eta^{3B_{d}}$ for all $h\in\mathcal{H}^{\prime}$. Let $\mathcal{X}$ be the set of $x\in[-CN^{d},CN^{d}]\cap\mathbb{Z}$ for which
$\min_{h^{\prime}\in\{0,h_{0}\}}\left\|\frac{x+h^{\prime}-a}{N^{\prime}}\right\|\geq\eta^{2B_{d}}.$
Then $|\mathcal{X}|\geq(2C-O(\eta^{2B_{d}}))N^{d}$. Note that, for any $x\in\mathcal{X}$, $h\in\mathcal{H}^{\prime\prime}$ and $m\in[\eta^{2B_{d}}N^{\prime}/2]$, we have
$\overline{\phi_{1,S}}(x+m+h)\phi_{1,S}(x+m)=1_{\alpha_{x}\not\in S}\cdot 1_{\alpha_{x+h}\not\in S}e((\alpha_{x}-\alpha_{x+h})(x+m)-h\alpha_{x+h}).$
Hence, recalling (4.19) and applying the pigeonhole principle, we see that there is some $x_{0}\in\mathcal{X}$ with $\alpha_{x_{0}}\not\in S$ such that for all $h\in\mathcal{H}^{\prime\prime}$ we have $\alpha_{x_{0}+h}\not\in S$ and
$|\mathbb{E}_{m\in[M^{\prime}]}e((\alpha_{x_{0}}-\alpha_{x_{0}+h})qm)|\gg_{C,P_{1},\ldots,P_{k}}\eta^{B_{d}}.$
By the geometric sum formula, we conclude that, for all $h\in\mathcal{H}^{\prime\prime}$, we have
$\|q(\alpha_{x_{0}}-\alpha_{x_{0}+h})\|\ll_{C,P_{1},\ldots,P_{k}}\frac{\eta^{-B_{d}}}{N^{\deg P_{1}}}.$
Recalling that $\alpha_{m}\in\frac{1}{N^{\deg P_{1}}}\mathbb{Z}$, by the pigeonhole principle we conclude that there is some constant $C_{d}^{\prime}\geq 1$ and some $\alpha\in\frac{1}{N^{\deg P_{1}}}\mathbb{Z}$ such that $\alpha_{m}=\alpha$ for $\geq\eta^{C_{d}^{\prime}}N^{d}$ integers $m\in[-CN^{d},CN^{d}]$. We have $\alpha\not\in S$, since $\alpha_{x_{0}+h}\not\in S$ for $h\in\mathcal{H}^{\prime\prime}$. Now the claim follows by writing $\phi_{1,S}=(\phi_{1}-\phi_{1,\{\alpha\}})+\phi_{1,S\cup\{\alpha\}}$ and applying the pigeonhole principle.
Step 3: Concluding the argument. Now, applying repeatedly the claim established above, starting with $S=\emptyset$ (in which case the assumption (4.16) holds for $\eta=\delta^{C_{d}^{\prime}}$ by (4.15)) and applying the above repeatedly, after $\ll\delta^{-C_{d}^{\prime}}$ iterations (4.18) cannot hold (since the number of different values that $\alpha_{m}$ takes at least $(\delta/2)^{C_{d}^{\prime}}N^{d}$ times is $\ll_{d}\delta^{-C_{d}^{\prime}}C$), so (4.17) holds with $\eta=\delta^{A_{d}}$ for some constant $A_{d}$.
Now that (4.17) holds with $\eta=\delta^{A_{d}}$, by the pigeonhole principle there exist intervals $[N_{1},N_{2}]\subset[1,N]$ of length $\delta^{3A_{d}C_{d}^{\prime}}N$ and $I\subset[-CN^{d},CN^{d}]$ of length $\delta^{3A_{d}C_{d}^{\prime}}N^{d}$ such that, denoting $\psi_{\alpha}(m)\coloneqq 1_{\alpha_{m}=\alpha}$, we have
$|\Lambda_{P_{1},\ldots,P_{k}}^{CN^{d},N}(\theta e(\alpha P_{1}(\cdot))1_{[N_{1},N_{2}]};\phi_{0}e(\alpha\cdot)1_{I},\psi_{\alpha},f_{3},\ldots,f_{k})|\geq\delta^{8A_{d}C_{d}^{\prime}}.$
But since the function $\psi_{\alpha}$ is constant on intervals of the form $[jN^{\prime}+a,(j+1)N^{\prime}+a)$ with $j\in\mathbb{Z}$, this implies
$|\Lambda_{P_{1},\ldots,P_{k}}^{CN^{d},N}(\theta e(\alpha P_{1}(\cdot))1_{[N_{1},N_{2}]};\phi_{0}e(\alpha\cdot)1_{I}\psi_{\alpha}(\cdot+P_{1}(N_{1})),f_{3},\ldots,f_{k})|\gg\delta^{8A_{d}C_{d}^{\prime}}.$
By the induction assumption and the fact that $d>\deg P_{1}$, for some $A_{d}^{\prime}$ we now obtain
$\|\theta 1_{[N_{1},N_{2}]}\|_{u^{d+1}[N]}\gg_{C,P_{1},\ldots,P_{k}}\delta^{A_{d}^{\prime}}.$
By Vinogradov’s Fourier expansion (Lemma 3.2), we can write
$1_{[N_{1},N_{2}]}(n)=\sum_{1\leq j\leq\delta^{-3A_{d}^{\prime}}}c_{j}e(\beta_{j}n)+E(n)$
for some real numbers $\beta_{j}$, some complex numbers $|c_{j}|\leq 1$ and some $|E(n)|\ll 1$ satisfying $\sum_{n\in[N]}|E(n)|\ll\delta^{2A_{d}^{\prime}}N$. Hence, we conclude that
$\|\theta\|_{u^{d+1}[N]}\gg_{C,P_{1},\ldots,P_{k}}\delta^{O_{C,P_{1},\ldots,P_{k}}(1)},$
as desired. ∎
4.4. Proof of Theorem 4.2
For the proof of Theorem 4.2, we need the following generalised von Neumann theorem for arithmetic progressions. This is well known (see for example [9, Lemma 2]), although the result is typically presented for functions defined on a cyclic group.
Lemma 4.6 (A generalised von Neumann theorem for arithmetic progressions).
Let $k\in\mathbb{N}$, $N\geq 1$, $C\geq 1$, and let $\theta\colon[N]\to\mathbb{C}$. Let $L_{1},\ldots,L_{k}$ be polynomials with integer coefficients of degree at most $1$ satisfying $L_{i}([N])\subset[-CN,CN]$ for all $i\in[k]$. Let $f_{1}\ldots,f_{k}\colon\mathbb{Z}\to\mathbb{C}$ be functions supported on $[-CN,CN]$ with $|f_{i}|\leq 1$ for $1\leq i\leq k$. Then we have
$\left|\frac{1}{N^{2}}\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\theta(n)f_{1}(m+L_{1}(n))\cdots f_{k}(m+L_{k}(n))\right|\ll_{k,C}\|\theta\|_{U^{k}[N]}.$
Lemma 4.6 could be proven directly without difficulty, but we deduce it as an immediate consequence of the following more general lemma that deals with multidimensional averages. This multidimensional version is needed for proving Theorem 1.5 (but for Theorem 4.2 the one-dimensional case suffices).
Lemma 4.7 (A multidimensional generalised von Neumann theorem for arithmetic progressions).
Let $k,r\in\mathbb{N}$, $N\geq 1$, $C\geq 1$, and let $\theta\colon[N]\to\mathbb{C}$. Let $L_{1},\ldots,L_{k}\colon\mathbb{Z}\to\mathbb{Z}^{r}$ be polynomials with integer coefficients of degree at most $1$ satisfying $L_{i}([N])\subset[-CN,CN]^{r}$ for all $i\in[k]$. Let $f_{1}\ldots,f_{k}\colon\mathbb{Z}^{r}\to\mathbb{C}$ be functions supported on $[-CN,CN]^{r}$ with $|f_{i}|\leq 1$ for $1\leq i\leq k$. Then we have
(4.20) $\begin{split}&\left|\frac{1}{N^{r+1}}\sum_{\mathbf{m}\in\mathbb{Z}^{r}}\sum_{n\in[N]}\theta(n)f_{1}(\mathbf{m}+L_{1}(n))\cdots f_{k}(\mathbf{m}+L_{k}(n))\right|\\ &\quad\ll_{k,r,C}\|\theta\|_{U^{k}[N]}.\end{split}$
Proof of Lemma 4.7.
For convenience, we extend the definition of $\theta$ to all of $\mathbb{Z}$ by setting it equal to 00 outside $[N]$.
We use induction on $k$. In the case $k=1$, the claim is immediate by making the change of variables $\mathbf{m}^{\prime}=\mathbf{m}+n\mathbf{v}_{1}+\mathbf{b}_{1}$ and noting that $\mathbb{E}_{n\in[N]}\theta(n)=\|\theta\|_{U^{1}[N]}$.
Suppose then that the claim holds in the case $k-1\in\mathbb{N}$ and consider the case $k$. Let $S$ be the expression inside the absolute values in (4.20). By making the change of variables $\mathbf{m}^{\prime}=\mathbf{m}+L_{1}(n)$, we have
$S=\frac{1}{N^{r+1}}\sum_{\mathbf{m}^{\prime}\in\mathbb{Z}^{r}}\sum_{n\in[N]}\theta(n)f_{1}(\mathbf{m}^{\prime})\prod_{j=2}^{k}f_{j}(\mathbf{m}^{\prime}+L_{j}(n)-L_{1}(n)).$
By the Cauchy–Schwarz inequality, we obtain
$|S|^{2}\ll_{r,C}\frac{1}{N^{r+2}}\sum_{\mathbf{m}\in\mathbb{Z}^{r}}\left|\sum_{n\in[N]}\theta(n)\prod_{j=2}^{k}f_{j}(\mathbf{m}+L_{j}(n)-L_{1}(n))\right|^{2}.$
In what follows, for a function $f\colon\mathbb{Z}^{r}\to\mathbb{C}$ and $\mathbf{v}\in\mathbb{Z}^{r}$, denote $\Delta_{h}^{\mathbf{v}}f(x)\coloneqq\overline{f(x+h\mathbf{v})}f(x)$. From van der Corput’s inequality (Lemma 3.1), we conclude that
$|S|^{2}\ll_{r,C}\frac{1}{N^{r+1}}\sum_{h\in\mathbb{Z}}\mu_{N}(h)\sum_{\mathbf{m}\in\mathbb{Z}^{r}}\sum_{n\in[N]}\Delta_{h}\theta(n)\prod_{j=2}^{k}\Delta_{h}^{\mathbf{v}_{j}}f_{j}(\mathbf{m}+L_{j}(n)-L_{1}(n)),$
where $\mathbf{v}_{j}\in\mathbb{Z}^{r}$ is such that $L_{j}(n)-L_{1}(n)=n\mathbf{v}_{j}+L_{j}(0)-L_{1}(0)$. By the induction assumption and the fact that $|\mu_{N}(h)|\ll\frac{1}{N}1_{|h|\leq N}$, we see that
(4.21) $|S|^{2}\ll_{k,r,C}\mathbb{E}_{h\in[-N,N]}\|\Delta_{h}\theta\|_{U^{k-1}[N]}.$
But by Hölder’s inequality and the definition of the $U^{k-1}[N]$ norm, the right-hand side of (4.21) is
$\ll_{k,C}\left(\mathbb{E}_{h\in[-N,N]}\mathbb{E}_{n,h_{1},\ldots,h_{k-1}\in[-N,N]}\prod_{\omega\in\{0,1\}^{k-1}}\mathcal{C}^{|\omega|}\Delta_{h}\theta(n+\omega\cdot(h_{1},\ldots,h_{k-1}))\right)^{1/2^{k-1}}$
$\ll_{k}\|\theta\|_{U^{k}[N]}^{2}.$
This completes the induction. ∎
Proof of Theorem 4.2.
For convenience, we extend the definition of $\theta$ to all of $\mathbb{Z}$ by setting it equal to 00 outside $[N]$.
We apply the PET induction scheme of Bergelson and Leibman [1]. For any finite collection $\mathcal{Q}$ of polynomials, define its type as $(d,w_{d},\ldots,w_{1})$, where $w_{j}$ is the number of different leading coefficients among the polynomials in the subcollection $\{Q\in\mathcal{Q}\colon\deg Q=j\}$. We introduce the lexicographic order $<$ on the types of polynomials. In other words, we write $(d,w_{d},\ldots,w_{1})<(d^{\prime},w_{d^{\prime}}^{\prime},\ldots,w_{1}^{\prime})$ if there exists $j\geq 0$ such that the first $j$ coordinates of the vectors are equal and the coordinate of order $j+1$ is larger for the second vector. In this way, we have introduced an order $<$ on the set of all finite collections of polynomials based on the order of their type. Note that the length of any descending chain of collections of polynomials with maximal element $\mathcal{Q}$ is bounded as a function of $|\mathcal{Q}|$ and $\deg\mathcal{Q}$.
We shall prove Theorem 4.2 by induction on the type of $\mathcal{Q}$. The base case is that of collections of type $(1,k)$ with $k\in\mathbb{N}$, that is, collections where all the polynomials have degree at most $1$. This case follows readily from Lemma 4.6.
Suppose that $\mathcal{Q}$ is a finite collection of polynomials for which Theorem 4.2 fails and that there is no smaller collection $\mathcal{Q}^{\prime}<\mathcal{Q}$ with this property. Then $\deg\mathcal{Q}\geq 2$. Let us write $\mathcal{Q}=\{Q_{1},\ldots,Q_{k}\}$, where $Q_{1}$ is one of the polynomials of $\mathcal{Q}$ with the least positive degree.
Let $S$ be the left-hand side of (4.2). By the Cauchy–Schwarz inequality, van der Corput’s inequality (3.6) and a change of variables, we have
(4.22) $\begin{split}|S|^{2}&\ll\frac{1}{N^{d+1}}\sum_{h\in\mathbb{Z}}\mu_{N}(h)\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\Delta_{h}\theta(n)\prod_{Q\in\mathcal{Q}}\Delta_{h}f_{Q}(m+Q(n))\\ &=\frac{1}{N^{d+1}}\sum_{h\in\mathbb{Z}}\mu_{N}(h)\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\Delta_{h}\theta(n)\prod_{\widetilde{Q}\in\mathcal{Q}_{h}}\widetilde{f}_{\widetilde{Q}}(m+\widetilde{Q}(n)),\end{split}$
where each $\widetilde{f}_{\widetilde{Q}}$ is equal to $f_{Q}$ or $\overline{f_{Q}}$ for some $Q\in\mathcal{Q}$ and $\mathcal{Q}_{h}$ is the collection (of size $\leq 2|\mathcal{Q}|$ and degree $\leq\deg\mathcal{Q}$) given by
$\mathcal{Q}_{h}=\{Q_{j}(y+h)-Q_{1}(y),Q_{j}(y)-Q_{1}(y)\}_{j=1}^{k}.$
We claim that for every $h\in\mathbb{Z}$ the type of $\mathcal{Q}_{h}$ is less than the type of $\mathcal{Q}$. Let $d_{1}=\deg Q_{1}$. Let $(d,w_{d},\ldots,w_{1})$ be the type of $\mathcal{Q}$ and let $(d_{h},w_{d,h},\ldots,w_{1,h})$ be the type of $\mathcal{Q}_{h}$. We have $d_{h}\leq d$, and if $d_{h}=d$, then $w_{e}=w_{e,h}$ for $d_{1}<e\leq d$. Moreover, $w_{d_{1},h}<w_{d_{1}}$, since if $c_{1},\ldots,c_{r}$ are the distinct leading coefficients of the degree $d_{1}$ polynomials in $\mathcal{Q}$, the leading coefficients of degree $d_{1}$ polynomials in $\mathcal{Q}_{h}$ are $c_{2}-c_{1},\ldots,c_{r}-c_{1}$. Hence we have $\mathcal{Q}_{h}<\mathcal{Q}$.
By (4.22), the estimate $|\mu_{N}(h)|\ll\frac{1}{N}1_{|h|\leq N}$ and the pigeonhole principle, we have
$|S|^{2}$ $\ll\mathbb{E}_{h\in[-N,N]}\max_{|j|\leq(C+1)N^{d-d_{h}}}\left|\frac{1}{N^{d_{h}+1}}\sum_{m\in\mathbb{Z}}1_{(jN^{d_{h}},(j+1)N^{d_{h}}]}(m)\sum_{n\in[N]}\Delta_{h}\theta(n)\prod_{\widetilde{Q}\in\mathcal{Q}_{h}}\widetilde{f}_{\widetilde{Q}}(m+\widetilde{Q}(n))\right|$
$=\mathbb{E}_{h\in[-N,N]}\max_{|j|\leq(C+1)N^{d-d_{h}}}\left|\frac{1}{N^{d_{h}+1}}\sum_{m^{\prime}\in\mathbb{Z}}\sum_{n\in[N]}\Delta_{h}\theta(n)1_{[N^{d_{h}}]}(m^{\prime})\prod_{\widetilde{Q}\in\mathcal{Q}_{h}}\widetilde{f}_{\widetilde{Q},jN^{d_{h}}}(m^{\prime}+\widetilde{Q}(n))\right|,$
where $\widetilde{f}_{\widetilde{Q},a}=\widetilde{f}_{\widetilde{Q}}(\cdot+a)$. Note that by the mean value theorem for all $\widetilde{Q}\in\mathcal{Q}_{h}$ we have $\max_{n\in[N]}|\widetilde{Q}(n)|\ll_{\deg\mathcal{Q}}N^{d_{h}}$, so for all $n\in[N]$ we have $\widetilde{Q}(n)\in[-C^{\prime}N^{d_{h}},C^{\prime}N^{d_{h}}]$ for some $C^{\prime}\ll_{\deg\mathcal{Q}}C$. Now, by the induction assumption (and the fact that the functions $\widetilde{f}_{Q,jN^{d}_{h}}$ may be assumed to be supported on $[-(C^{\prime}+1)N^{d_{h}},(C^{\prime}+1)N^{d_{h}}]$), we conclude that for some natural number $s^{\prime}\ll_{|\mathcal{Q}|,\deg{Q}}$ we have
$|S|^{2}\ll_{|\mathcal{Q}|,\deg\mathcal{Q},C}\mathbb{E}_{h\in[-N,N]}\|\Delta_{h}\theta\|_{U^{s^{\prime}}[N]}.$
Applying Hölder’s inequality as in the proof of Lemma 4.6, this implies that
$|S|\ll_{|\mathcal{Q}|,\deg{Q},C}\|\theta\|_{U^{s^{\prime}+1}[N]}.$
This completes the induction. ∎
5. Lemmas for the main proofs
5.1. Quantitative uniformity bounds for the Möbius function
In this subsection, we state the quantitative bounds for the uniformity of the Möbius function that we need for the proofs of the main theorems.
Lemma 5.1 (Quantitative $U^{k}$- and $u^{k}$-uniformity of $\mu$).
Let $k\in\mathbb{N}$, $A>0$ and $N\geq 3$.
  1. (i)
    We have
    $\|\mu\|_{U^{4}[N]}\ll_{A}(\log N)^{-A}.$
  2. (ii)
    We have
    $\|\mu\|_{U^{k}[N]}\ll_{k}(\log\log N)^{-c_{k}}$
    for some $c_{k}>0$.
  3. (iii)
    We have
    $\|\mu\|_{u^{k}[N]}\ll_{A,k}(\log N)^{-A}.$
Proof.
Part (i) follows by combining recent work of Leng [22, Theorems 4 and 5] with the triangle inequality for the Gowers norms. Part (ii) was proven in [26, Theorem 1.4]. Recalling Definition 3.5 of the $u^{k}[N]$ norm, part (iii) follows from the Green–Tao Möbius and nilsequences estimate [13, Theorem 1.1] (but it would also be possible to prove it by more classical means by using estimates for bilinear sums of polynomial exponential phases). ∎
Remark 5.2.
Prior to [22], Leng proved in [21] that $\|\mu\|_{U^{3}[N]}\ll_{A}(\log N)^{-A}$. This result would already suffice for proving Theorem 1.2(i) for $k=2$; see Theorem  6.1.
Remark 5.3.
Lemma 5.1 continues to hold if the Möbius function $\mu$ is replaced with the Liouville function $\lambda$. This follows easily from the identities
$\lambda(n)=\sum_{d\geq 1}\mu(n/d)1_{d^{2}\mid n},\quad\textnormal{and}\quad\mu(n/d)1_{d^{2}\mid n}=\sum_{\begin{subarray}{c}n=en^{\prime}\\ d^{2}\mid e\mid d^{\infty}\\ (d,n^{\prime})=1\end{subarray}}\mu(e/d)\mu(n^{\prime}),$
which can be truncated to $d^{2}\leq e\leq K$ for any $K\geq 1$ at the cost of an error term that is bounded by $O(1/K)$ in $L^{1}[N]$ norm.
5.2. A lacunary subsequence trick
We use a lacunary subsequence trick in the proofs of our pointwise convergence results. Such a trick roughly states that if $A_{N}\colon X\to\mathbb{C}$ are some measurable functions and we have strong quantitative decay for $\|A_{N_{j}}\|_{L^{1}(X)}$ for some lacunary sequence $(N_{j})_{j\in\mathbb{N}}$, then provided that $A_{N}$ does not vary too much on intervals of the form $[N_{j},N_{j+1}]$ in $L^{1}(X)$ norm, the sequence $A_{N}$ must converge to 00 in $L^{\infty}(X)$ norm. Variants of this idea are frequently used to establish convergence of ergodic averages; see for example [10, Section 5].
Lemma 5.4.
Let $B>1$ and $c>0$. Let $(X,\nu)$ be a probability space, and for $N\geq 1$ let $A_{N}\colon X\to\mathbb{C}$ be a measurable function.
  1. (i)
    If for any $N\geq 3$ we have
    (5.1) $\|A_{N}\|_{L^{1}(X)}\ll(\log N)^{-B}$
    and
    (5.2) $\left\|\sup_{M\in[N,(1+(\log N)^{-(B/2-1)})N]}|A_{M}-A_{N}|\right\|_{L^{1}(X)}\ll(\log N)^{-B/3},$
    then, for any $\varepsilon>0$, we have $\lim_{N\to\infty}(\log N)^{B/4}\|A_{N}\|_{L^{\infty}(X)}=0$.
  2. (ii)
    If for any $N\geq 3$ we have
    (5.3) $\|A_{N}\|_{L^{1}(X)}\ll(\log\log N)^{-c}$
    and
    (5.4) $\begin{split}\left\|\sup_{M\in[N,\exp((\log N)^{1+(\log\log N)^{-c}})]}|A_{M}-A_{N}|\right\|_{L^{1}(X)}\ll(\log\log N)^{-c},\end{split}$
    then we have $\lim_{N\to\infty}(\log\log N)^{c/3}\|A_{N}\|_{L^{\infty}(X)}=0$.
  3. (iii)
    If for any $N\geq 3$ we have
    (5.5) $\|A_{N}\|_{L^{1}(X)}\ll(\log\log N)^{-c},$
    then, for any sequence $(N_{j})_{j\in\mathbb{N}}$ satisfying $N_{j}\gg\exp(\exp(j^{2/c}))$, we have
    $\lim_{j\to\infty}(\log\log N_{j})^{c/3}\|A_{N_{j}}\|_{L^{\infty}(X)}=0$.
Proof.
(i) Let $\eta>0$ be a small enough constant, and set $N_{j}=\exp(\eta\cdot j^{2/B})$ for all $j\in\mathbb{N}$. Note that $N_{j+1}\leq N_{j}(1+(\log N_{j})^{-(B/2-1)})$ for all large enough $j\in\mathbb{N}$. For $\delta>0$, denote
(5.6) $\begin{split}E(\delta)&=\{x\in X\colon\limsup_{N\to\infty}\,(\log N)^{B/4}|A_{N}(x)|\geq\delta\},\\ E_{j}(\delta)&=\{x\in X\colon|A_{N_{j}}(x)|\geq\frac{\delta}{3}(\log N_{j})^{B/4}\},\\ E_{j}^{\prime}(\delta)&=\{x\in X\colon\sup_{M\in[N_{j},N_{j+1}]}|A_{M}(x)-A_{N}(x)|\geq\frac{\delta}{3}(\log N_{j})^{B/4}\}.\end{split}$
Then we have
(5.7) $E(\delta)\subset\bigcap_{i=3}^{\infty}\bigcup_{j\geq i}(E_{j}(\delta)\cup E_{j}^{\prime}(\delta)).$
By Markov’s inequality, (5.1) and (5.2), we have the bounds
$\nu(E_{j}(\delta))$ $\ll_{\delta}(\log N_{j})^{B/4-B}\ll j^{-3/2},$
$\nu(E_{j}^{\prime}(\delta))$ $\ll_{\delta}(\log N_{j})^{B/4-B}\ll j^{-3/2}.$
Since $\sum_{j\geq 3}j^{-3/2}<\infty$, from (5.7) and the Borel–Cantelli lemma we conclude that $\nu(E(\delta))=0$. By the countable additivity of $\nu$, this then implies that
$\nu(\bigcap_{\delta>0}E(\delta))=\nu(\bigcap_{j\geq 1}E(1/j))=0,$
as desired.
(ii) Let $\eta>0$ be a small enough constant, and set $N_{j}=\exp(\exp(\eta\cdot j^{2/c}))$ for all $j\in\mathbb{N}$. Then we have $N_{j+1}\leq\exp((\log N_{j})^{1+(\log\log N_{j})^{-c/2}})$ for all large enough $j\in\mathbb{N}$. For $\delta>0$, denote
$E(\delta)$ =$=\{x\in X\colon\limsup_{N\to\infty}\,(\log\log N)^{c/3}|A_{N}(x)|\geq\delta\}$
$E_{j}(\delta)$ =$=\{x\in X\colon|A_{N_{j}}(x)|\geq\frac{\delta}{3}(\log\log N_{j})^{-c/3}\},$
$E_{j}^{\prime}(\delta)$ =$=\{x\in X\colon\sup_{M\in[N_{j},N_{j+1}]}|A_{M}(x)-A_{N}(x)|\geq\frac{\delta}{3}(\log\log N_{j})^{-c/3}\}.$
Then we have (5.7). From (5.3), (5.4) we have
$\nu(E_{j}(\delta)),\nu(E_{j}^{\prime}(\delta))\ll_{\delta}(\log\log N_{j})^{-2c/3}\ll j^{-4/3}$
for all $j\geq 3$, so as above the Borel–Cantelli lemma gives $\nu(\bigcap_{\delta>0}E(\delta))$, which was the claim.
(iii) Let a sequence $N_{j}\gg\exp(\exp(j^{2/c}))$ be given. For $\delta>0$, denote
$E(\delta)$ =$=\{x\in X\colon\limsup_{i\to\infty}\,(\log\log N_{i})^{c/3}|A_{N_{i}}(x)|\geq\delta\},$
$E_{j}(\delta)$ =$=\{x\in X\colon|A_{N_{j}}(x)|\geq\frac{\delta}{2}(\log\log N_{j})^{-c/3}\}.$
Then we have
$E(\delta)\subset\bigcap_{i=1}^{\infty}\bigcup_{j\geq i}E_{j}(\delta).$
By Markov’s inequality and (5.5), we have
$\nu(E_{j}(\delta))\ll_{\delta}(\log\log N_{j})^{-2c/3}\ll j^{-4/3}.$
Since $\sum_{j\geq 2}j^{-4/3}<\infty$, the Borel–Cantelli lemma again gives us $\nu(\bigcap_{\delta>0}E(\delta))=0$, as desired. ∎
5.3. A simple $L^{1}$ bound
We will also need a simple $L^{1}$ estimate for ergodic averages that follows from Hölder’s inequality.
Lemma 5.5.
Let $k\in\mathbb{N}$. Let $(X,\nu)$ be a probability space, and let $T_{1},\ldots,T_{k}\colon X\to X$ be invertible measure-preserving maps. Let $g_{1},\ldots,g_{k}\colon\mathbb{Z}\to\mathbb{Z}$ be functions. Let $S\subset\mathbb{N}$ be a finite and nonempty. Also let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}\leq 1$. Then, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$, we have
$\|\mathbb{E}_{n\in\mathcal{S}}f_{1}(T_{1}^{g_{1}(n)}x)\cdots f_{k}(T_{k}^{g_{k}(n)}x)\|_{L^{1}(X)}\leq\|f_{1}\|_{L^{q_{1}}(X)}\cdots\|f_{k}\|_{L^{q_{k}}(X)}.$
Proof.
Let $r\geq 1$ be such that $\frac{1}{r}=\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}$. Then, by Jensen’s and Hölder’s inequalities and the $T$-invariance of $\nu$, we have
$\|\mathbb{E}_{n\in\mathcal{S}}f_{1}(T_{1}^{g_{1}(n)}x)\cdots f_{k}(T_{k}^{g_{k}(n)}x)\|_{L^{1}(X)}$
$\quad\leq\|\mathbb{E}_{n\in\mathcal{S}}f_{1}(T_{1}^{g_{1}(n)}x)\cdots f_{k}(T_{k}^{g_{k}(n)}x)\|_{L^{r}(X)}$
$\quad\leq\left(\int_{X}\mathbb{E}_{n\in\mathcal{S}}|f_{1}(T_{1}^{g_{1}(n)}x)|^{q_{1}}\textnormal{ d}\nu(x)\right)^{1/q_{1}}\cdots\left(\int_{X}\mathbb{E}_{n\in\mathcal{S}}|f_{k}(T_{k}^{g_{k}(n)}x)|^{q_{k}}\textnormal{ d}\nu(x)\right)^{1/q_{k}}$
$\quad=\|f_{1}\|_{L^{q_{1}}(X)}\cdots\|f_{k}\|_{L^{q_{k}}(X)},$
as claimed. ∎
6. Proofs of the pointwise ergodic theorems
All of our main theorems will be proven in the more general setting of weighted polynomial ergodic averages
(6.1) $A_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1},\ldots,f_{k})(x)\coloneqq\frac{1}{N}\sum_{n\leq N}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x),$
where $\theta\colon\mathbb{N}\to\mathbb{C}$ is a function satisfying suitable hypotheses. Some of these hypotheses turn out to be trivial when $|\theta|\leq 1$ (so in particular for $\theta=\mu$), but the hypotheses are tailored so that the results are also applicable to a suitably normalised version of the von Mangoldt function (by the main results of [22][26]).
We start with Theorem 1.2. This theorem is an immediate consequence of Lemma 5.1(i),(iii) and the following theorem concerning the averages (6.1).
Theorem 6.1 (Pointwise convergence of some polynomial ergodic averages with nice weight).
Let $d,k,K\in\mathbb{N}$. Let $\theta\colon\mathbb{N}\to\mathbb{C}$ satisfy $|\theta(n)|\leq(\log n)^{K}$ for all $n\in\mathbb{N}$. Suppose also that, for any $M\geq 3$ and $B>0$, we have
(6.2) $\begin{cases}\|\theta\|_{U^{k+1}[M]}\ll_{B}(\log M)^{-B}\quad&\textnormal{if}\quad d=1\\ \|\theta\|_{u^{d+1}[M]}\ll_{B}(\log M)^{-B}\quad&\textnormal{if}\quad d\geq 2.\end{cases}$
Let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients satisfying $\deg P_{1}\leq\cdots\leq\deg P_{k}=d$. Suppose that one of the following holds:
  1. (i)
    $d=1;$
  2. (ii)
    $d\geq 2$ and $P_{1},\ldots,P_{k}$ have pairwise distinct degrees.
Let $(X,\nu,T)$ be a measure-preserving system, and let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$. Then, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$ and $A\geq 0$, we have
$\lim_{N\to\infty}\frac{(\log N)^{A}}{N}\sum_{n\leq N}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)=0$
for almost all $x\in X$.
Remark 6.2.
We briefly comment on what Theorem 6.1 tells us about the convergence of
(6.3) $\mathbb{E}_{n\in[N]}\theta(n)f_{1}(T^{n}x)f_{2}(T^{P(n)}x)$
for other bounded multiplicative functions $\theta$ than the Möbius function. Firstly, if we aim for convergence to 00, we should assume that $\theta$ is aperiodic (that is, it has mean 00 on every arithmetic progression). If we make no further assumption on $\theta$ than aperiodicity, even the $U^{1}[N]$ norm of $\theta$ could tend to zero arbitrarily slowly, so Theorem 6.1 cannot be applied in this case. But if we assume more strongly that $\theta$ satisfies the Siegel–Walfisz assumption (that is, $|\sum_{n\in\mathcal{P}}\theta(n)|\ll_{A}N/(\log N)^{A}$ for any arithmetic progression $\mathcal{P}\subset[N]$), and that $\theta$ has a combinatorial decomposition into type I and II sums, then it should be possible to verify the second bound in (6.2), leading to the convergence of (6.3) in the case $\deg P\geq 2$. But to obtain the first bound in (6.2) (which corresponds to $\deg P=1$ in (6.3)) we would need to impose an assumption stronger than the Siegel–Walfisz assumption in view of the best known $U^{3}$ inverse theorem [21, Theorem 7], which is quasipolynomial.
We shall reduce Theorem 6.1 to the following quantitative $L^{1}$ estimate.
Proposition 6.3.
Let the notation and assumptions be as in Theorem 6.1. Then, for all $N\geq 3$, we have
(6.4) $\|A_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1},\ldots,f_{k})\|_{L^{1}(X)}\ll_{A,K,P_{1},\ldots,P_{k}}\left(\prod_{j=1}^{k}(1+\|f_{j}\|_{L^{q_{j}}(X)})\right)^{B(q_{1},\ldots,q_{k})}(\log N)^{-A}$
for some $B(q_{1},\ldots,q_{k})>0$.
Proof that Proposition 6.3 implies Theorem 6.1.
We are going to apply Lemma 5.4(i). Recall the notation (6.1). By Lemma 5.5 and the assumption $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$, for any $B>0$, $N\geq 3$ and $M\in[N,(1+(\log N)^{-B})N]$, we have
$(\log N)^{-K}\|A_{M}^{P_{1},\ldots,P_{k}}(\theta;f_{1},\ldots,f_{k})-A_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1},\ldots,f_{k})\|_{L^{1}(X)}$
$\quad\leq\int_{X}\left(\frac{1}{N}\sum_{M\leq n\leq N}\prod_{j=1}^{k}|f_{j}(T^{P_{j}(n)}x)|+\left(\frac{1}{M}-\frac{1}{N}\right)\sum_{n\leq N}\prod_{j=1}^{k}|f_{j}(T^{P_{j}(n)}x)|\right)\textnormal{ d}\nu(x)$
$\quad\ll(\log N)^{-B}\|f_{1}\|_{L^{q_{1}}(X)}\cdots\|f_{k}\|_{L^{q_{k}}(X)}.$
Hence, by Lemma 5.4(i), the conclusion of Theorem 6.1 follows from (6.4). ∎
Proof of Proposition 6.3.
We first reduce the proof of Proposition 6.3 to the case $q_{1}=\cdots=q_{k}=\infty$.
Reduction to the case of bounded functions. We claim that it suffices to prove Proposition 6.3 in the case $q_{1}=\cdots=q_{k}=\infty$. Suppose that this case has been proven. Let $1<q_{1},\ldots,q_{k}<\infty$ and $f_{i}\in L^{q_{i}}(X)$ for $i\in\{1,\ldots,k\}$ (we can assume that all the $q_{i}$ are <absent<$\infty$< ∞, since we have $\|f\|_{L^{q}(X)}\leq\|f\|_{L^{\infty}(X)}$ for any $q<\infty$ and any measurable $f$). Also let $A>0$ and let $C>0$ be large enough in terms of $A,K,q_{1},\ldots q_{k}$.
For $j\in\{1,\ldots,k\}$ and $N\geq 2$, we split
$f_{j}=f_{j,N}^{(1)}+f_{j,N}^{(2)},\quad\textnormal{where}\quad f_{j,N}^{(1)}(x)=f_{j}(x)1_{|f_{j}(x)|\leq(\log N)^{C}},\,\,f_{j,N}^{(2)}(x)=f_{j}(x)1_{|f_{j}(x)|>(\log N)^{C}}.$
Then by linearity we see that
(6.5) $\begin{split}&A_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1},\ldots,f_{k})\\ \quad&=A_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1,N}^{(1)},\ldots,f_{k,N}^{(1)})+\sum_{(i_{1},\ldots,i_{k})\in[2]^{k}\setminus(1,\ldots,1)}A_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1,N}^{(i_{1})},\ldots,f_{k,N}^{(i_{k})}).\end{split}$
Since $\|(\log N)^{-C}f_{j,N}^{(1)}\|_{L^{\infty}(X)}\leq 1$ for $j\in\{1,\ldots,k\}$, by the case $q_{1}=\cdots=q_{k}=\infty$ we have
$\|A_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1,N}^{(1)},\ldots,f_{k,N}^{(1)})\|_{L^{1}(X)}\ll_{A,C,K,P_{1},\ldots,P_{k}}(\log N)^{-A}.$
To bound the error term in (6.5), it suffices to show that for $(i_{1},\ldots,i_{k})\in[2]^{k}\setminus(1,\ldots,1)$ we have
$\|A_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1,N}^{(i_{1})},\ldots,f_{k,N}^{(i_{k})})\|_{L^{1}(X)}\ll_{C,q_{1},q_{2}}(\log N)^{K-cC}\|f_{1}\|_{L^{q_{1}}(X)}\cdots\|f_{k}\|_{L^{q_{k}}(X)}$
for some $c=c(q_{1},\ldots,q_{k})>0$, since $C$ can be taken to be large enough in terms of $A,K,q_{1},\ldots,q_{k}$ so that $K-cC\leq-A$.
Fix some $q_{i}^{\prime}\in(1,q_{i})$ such that $\frac{1}{q_{1}^{\prime}}+\cdots+\frac{1}{q_{k}^{\prime}}=1$, and let $c=\frac{q_{1}}{q_{1}^{\prime}}+\cdots+\frac{q_{k}}{q_{k}^{\prime}}-k$. Recalling the assumption $|\theta(n)|\leq(\log n)^{K}$ and using Lemma 5.5, we get
$\|A_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1,N}^{(i_{1})},\ldots,f_{k,N}^{(i_{k})})\|_{L^{1}(X)}$
$\quad\leq(\log N)^{K-cC}\|A_{N}^{P_{1},\ldots,P_{k}}(1;|f_{1,N}^{(i_{1})}|^{q_{1}/q_{1}^{\prime}},\ldots,|f_{k,N}^{(i_{k})}|^{q_{k}/q_{k}^{\prime}})\|_{L^{1}(X)}$
$\quad\ll(\log N)^{K-cC}\|f_{1}\|_{L^{q_{1}}(X)}^{q_{1}/q_{1}^{\prime}}\cdots\|f_{k}\|_{L^{q_{k}}(X)}^{q_{k}/q_{k}^{\prime}}.$
Now the proof of Proposition 6.3 has been reduced to the case $q_{1}=\cdots=q_{k}=\infty$.
The case of bounded functions. It now remains to show (6.4) in the case $q_{1}=\cdots=q_{k}=\infty$. In the rest of the proof, we abbreviate $A_{N}(x)\coloneqq A_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1},\ldots,f_{k})(x)$.
By Cauchy–Schwarz, it suffices to show that
$\int_{X}|A_{N}(x)|^{2}\textnormal{ d}\nu(x)\ll_{A,K,a,P}\|f_{1}\|_{L^{\infty}(X)}^{2}\cdots\|f_{k}\|_{L^{\infty}(X)}^{2}(\log N)^{-A}.$
The assumption $|\theta(n)|\leq(\log n)^{K}$ gives $\|A_{N}\|_{L^{\infty}(X)}\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}(\log N)^{K}$, so it suffices to show that for any $\phi\in L^{\infty}(X)$ we have
(6.6) $\int_{X}\phi(x)A_{N}(x)\textnormal{ d}\nu(x)\ll_{A,K,a,P}\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log N)^{-A}.$
By the $T$-invariance of $\nu$, the claim (6.6) is equivalent to
(6.7) $\begin{split}&\int_{X}\frac{1}{N^{d+1}}\sum_{m\in[N^{d}],n\in[N]}\theta(n)\phi(T^{m}x)f_{1}(T^{m+P_{1}(n)}x)\cdots f_{k}(T^{m+P_{k}(n)}x)\textnormal{ d}\nu(x)\\ &\quad\ll_{A,K,P_{1},\ldots,P_{k}}\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log N)^{-A}.\end{split}$
By the definition of the $L^{\infty}(X)$ norm, there exists a set $X^{\prime}\subset X$ such that
$\nu(X^{\prime})=1\quad\textnormal{and}\quad|\phi(x)|\leq\|\phi\|_{L^{\infty}(X)},\,|f_{i}(x)|\leq\|f_{i}\|_{L^{\infty}(X)}\,\,\textnormal{for all}\,\,1\leq i\leq k,\,x\in X^{\prime}.$
Restricting the integral in (6.7) to $X^{\prime}$, it suffices to show that for all $x\in X^{\prime}$ we have the bound
(6.8) $\begin{split}&\left|\frac{1}{N^{d+1}}\sum_{m\in[N^{d}],n\in[N]}\frac{\theta(n)}{(\log N)^{K}}\phi(T^{m}x)f_{1}(T^{m+P_{1}(n)}x)\cdots f_{k}(T^{m+P_{k}(n)}x)\right|\\ &\quad\ll_{A,P_{1},\ldots,P_{k}}\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log N)^{-2-K-A}.\end{split}$
For proving this estimate, we may assume that $\phi,f_{1},\ldots,f_{k}$ are supported on $[-CN^{,}CN^{d}]$ for some $C\ll_{a,P}1$. Note also that by assumption for $n\in[N]$ we have $|\theta(n)|/(\log N)^{K}\leq 1$.
If $d=1$ (that is, all the $P_{i}$ of degree at most $1$), the estimate (6.8) follows by combining Lemma 4.6 and the assumption (6.2). If instead $d\geq 2$ (in which case by assumption the $P_{i}$ have distinct degrees), (6.8) follows by combining Theorem 4.1 and the assumption (6.2). ∎
We then turn to Theorem 1.4. This theorem is an immediate consequence of the following general result combined with Lemma 5.1(ii).
Theorem 6.4 (Pointwise convergence of double-logarithmic polynomial ergodic averages with nice weight).
Let $K\in\mathbb{N}$. Let $\theta\colon\mathbb{N}\to\mathbb{C}$ satisfy $|\theta(n)|\leq K\log n$ for all $n\in\mathbb{N}$, and suppose that there exist some infinite sets $S_{1},\ldots,S_{K}\subset\mathbb{N}$ such that, for all large enough $N\geq 1$ and any $n\in[N]$, we have
(6.9) $|\theta(n)|\leq\sum_{j=1}^{K}\frac{N}{|S_{j}\cap[N]|}1_{S_{j}}(n).$
Suppose also that for all $s\in\mathbb{N}$ and $M\geq 3$ we have
(6.10) $\|\theta\|_{U^{s}[M]}\ll_{s}(\log\log M)^{-\gamma_{s}}$
for some $\gamma_{s}>0$.
Let $(X,\nu,T)$ be a measure-preserving system. Let $k\in\mathbb{N}$, let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients, and let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$. Then, there exists $c=c(P_{1},\ldots,P_{k},q_{1},\ldots,q_{k})>0$ such that, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$ and $N\geq 3$, we have
(6.11) $\lim_{N\to\infty}\,(\log\log N)^{c}\mathbb{E}_{n\leq N}^{\log\log}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)=0$
for almost all $x\in X$.
Theorem 6.1 will also be reduced to an $L^{1}$ estimate.
Proposition 6.5.
Let the notation and assumptions be as in Theorem 6.4. Then, for all $N\geq 3$, we have
$\|\mathbb{E}_{n\leq N}^{\log\log}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)\|_{L^{1}(X)}$
$\quad\ll_{K,P_{1},\ldots,P_{k}}((1+\|f_{1}\|_{L^{q_{1}}(X)})\cdots(1+\|f_{k}\|_{L^{q_{k}}(X)}))^{B(q_{1},\ldots,q_{k})}(\log\log N)^{-c^{\prime}}$
for some $c=c(P_{1},\ldots,P_{k})>0$ and $B(q_{1},\ldots,q_{k})>0$.
Proof of Theorem 6.4 assuming Proposition 6.5.
We are going to apply Lemma 5.4(ii). Note that for any bounded sequence $(a(n))_{n\in\mathbb{N}}$ and any $M\geq N$ we have
$\left|\frac{1}{\log\log M}\sum_{n\leq M}\frac{a(n)}{n\log n}-\frac{1}{\log\log N}\sum_{n\leq N}\frac{a(n)}{n\log n}\right|$
$\quad\ll\left(\frac{1}{\log\log N}-\frac{1}{\log\log M}\right)\sum_{n\leq M}\frac{|a(n)|}{n\log n}+\frac{1}{\log\log M}\sum_{N\leq n\leq M}\frac{|a(n)|}{n\log n}$
$\quad\ll\frac{\log\log M-\log\log N+1}{\log\log N}\frac{1}{\log\log M}\sum_{\ell\leq(\log M)/(\log 2)+1}\frac{1}{\ell}\mathbb{E}_{n\leq 2^{\ell}}|a(n)|$
$\quad\quad+\frac{1}{\log\log M}\sum_{(\log N)/(\log 2)\leq\ell\leq(\log M)/(\log 2)+1}\frac{1}{\ell}\mathbb{E}_{n\leq 2^{\ell}}|a(n)|.$
Using this with $a(n)=\theta(n)\prod_{j\in[k]}f_{j}(T^{P_{j}(n)}x)$, $M\in[N,\exp((\log N)^{1+(\log\log N)^{-c}})]$, and using Lemma 5.5 and the assumption (6.9), we obtain
$\|\mathbb{E}_{n\leq M}^{\log\log}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)-\mathbb{E}_{n\leq N}^{\log\log}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)\|_{L^{1}(X)}$
$\quad\ll(\log\log N)^{-c}\max_{j\in[K]}\max_{\ell\leq(\log M)/(\log 2)+1}\left(\mathbb{E}_{n\in S_{j}\cap[2^{\ell}]}\int_{X}|f_{1}(T^{P_{1}(n)}x)|\cdots|f_{k}(T^{P_{k}(n)}x)|\textnormal{ d}\nu(x)\right)$
$\quad\ll(\log\log N)^{-c}\|f_{1}\|_{L^{q_{1}}(X)}\cdots\|f_{k}\|_{L^{q_{k}}(X)}.$
Hence, from Lemma 5.4(ii) it follows that Proposition 6.5 implies Theorem 1.2. ∎
Proof of Proposition 6.5.
We again first reduce to the bounded case.
Reduction to the case of bounded functions. Suppose that the case $q_{1}=\ldots=q_{k}=\infty$ has been proven. Let $1<q_{1},\ldots,q_{k}<\infty$, $f_{i}\in L^{q_{i}}(X)$, and let $\kappa>0$ be small enough in terms of $P_{1},\ldots,P_{k}$. For $i\in\{1,\ldots,k\}$, we split
$f_{i}=f_{i,N}^{(1)}+f_{i,N}^{(2)},\quad\textnormal{where}\quad f_{i,N}^{(1)}(x)=f_{i}(x)1_{|f_{i}(x)|\leq(\log\log N)^{\kappa}},\,\,f_{i,N}^{(2)}(x)=f_{i}(x)1_{|f_{i}(x)|>(\log\log N)^{\kappa}}.$
Then, denoting by $\widetilde{A}_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1},\ldots,f_{k})$ the function inside the norm in (6.11), by linearity we obtain
$\widetilde{A}_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1},\ldots,f_{k})$ $=\widetilde{A}_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1,N}^{(1)},\ldots,f_{k,N}^{(1)})$
$\quad+\sum_{(i_{1},\ldots,i_{k})\in[2]^{k}\setminus(1,\ldots,1)}\widetilde{A}_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1,N}^{(i_{1})},\ldots,f_{k,N}^{(i_{k})}).$
Recalling that $\kappa>0$ is small enough, by the case of bounded functions we have
$\|\widetilde{A}_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1,N}^{(1)},\ldots,f_{k,N}^{(1)})\|_{L^{1}(X)}\ll_{k,P_{1},\ldots,P_{k},\kappa}(\log\log N)^{-c/2},$
where $c=c(P_{1},\ldots,P_{k})$ is as in (6.11) in the case of bounded functions.
Note that for any $q^{\prime}<q$ and $i\in\{1,\ldots,k\}$ we have
(6.12) $\int_{X}|f_{i,N}^{(2)}(x)|^{q^{\prime}}\textnormal{ d}\nu(x)\leq(\log\log N)^{-\kappa\cdot(q^{\prime}-q)}\int_{X}|f_{i}(x)|^{q}\textnormal{ d}\nu(x).$
Fix some $q_{i}^{\prime}\in(1,q_{i})$ with $\frac{1}{q_{1}^{\prime}}+\ldots+\frac{1}{q_{k}^{\prime}}=1$. Note also that, by the assumption on $\theta$, for $n\in[N]$ we have
$|\theta(n)|\leq\theta_{1}(n)+\cdots+\theta_{K}(n),\quad\textnormal{where}\quad\theta_{j}(n)=\frac{N}{|S_{j}\cap[N]|}1_{S_{j}}(n).$
Now, by Lemma 5.5 and (6.12), for any $(i_{1},\ldots,i_{k})\neq(1,\ldots,1)$ we get
$\|\widetilde{A}_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1,N}^{(i_{1})},\ldots,f_{k,N}^{(i_{k})})\|_{L^{1}(X)}$
$\quad\ll_{K}\max_{1\leq j\leq J}\|\widetilde{A}_{N}^{P_{1},\ldots,P_{k}}(\theta_{j};f_{1,N}^{(i_{1})},\ldots,f_{k,N}^{(i_{k})})\|_{L^{1}(X)}$
$\quad\leq(\log\log N)^{-c^{\prime}\kappa}\|f_{1}\|_{L^{q_{1}}(X)}\cdots\|f_{k}\|_{L^{q_{k}}(X)},$
where $c^{\prime}=\frac{q_{1}}{q_{1}^{\prime}}+\cdots+\frac{q_{k}}{q_{k}^{\prime}}-k>0$.
The case of bounded functions. Let $P_{1},\ldots,P_{k}$ be fixed polynomials with integer coefficients and with highest degree $d\in\mathbb{N}$. Let $(X,\nu,T)$ be a measure-preserving system, and let $f_{1},\ldots,f_{k}\in L^{\infty}(X)$. In the rest of the proof, we abbreviate $\widetilde{A}_{N}(x)\coloneqq\widetilde{A}_{N}(\theta;f_{1},\ldots,f_{k})(x)$.
By Cauchy–Schwarz, it suffices to show that
$\int_{X}|\widetilde{A}_{N}(x)|^{2}\textnormal{ d}\nu(x)\ll\|f_{1}\|_{L^{\infty}(X)}^{2}\cdots\|f_{k}\|_{L^{\infty}(X)}^{2}(\log\log N)^{-c}.$
From (6.9) we have $\limsup_{N\to\infty}\|\widetilde{A}_{N}\|_{L^{\infty}(X)}\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}$, so it suffices to show that for any $\phi\in L^{\infty}(X)$ we have
$\int_{X}\phi(x)\widetilde{A}_{N}(x)\textnormal{ d}\nu(x)\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N)^{-c}.$
This can be written out as
$\int_{X}\mathbb{E}_{n\leq N}^{\log\log}\theta(n)\phi(x)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)\textnormal{ d}\nu(x)$
$\quad\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N)^{-c}.$
Let $N_{0}=\exp(\exp((\log\log N)^{1/2}))$. By partial summation and the assumptions (6.9) and $|\theta(n)|\leq K\log n$, we have the bound
$\sum_{n\leq N_{0}}\frac{|\theta(n)|}{n\log n}\ll_{K}(\log\log N)^{1/2},$
so it suffices to show that
$\int_{X}\frac{1}{\log\log N}\sum_{N_{0}\leq n\leq N}\frac{\theta(n)}{n\log n}\phi(x)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)\textnormal{d}\nu(x)$
$\quad\ll(\log\log N)^{-c}\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}.$
By partial summation and the assumptions on $\theta$, this follows if we show uniformly for $N^{\prime}\in[N_{0},N]$ the bound
$\int_{X}\frac{1}{N^{\prime}}\sum_{n\in[N^{\prime}]}\theta(n)\phi(x)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)\textnormal{ d}\nu(x)$
$\quad\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N)^{-c}.$
By the $T$-invariance of $\nu$, it suffices to show that
(6.13) $\begin{split}&\int_{X}\frac{1}{(N^{\prime})^{d+1}}\sum_{m\in[(N^{\prime})^{d}],n\in[N^{\prime}]}\theta(n)\phi(T^{m}x)f_{1}(T^{m+P_{1}(n)}x)\cdots f_{k}(T^{m+P_{k}(n)}x)\textnormal{ d}\nu(x)\\ &\quad\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N)^{-c}.\end{split}$
We can restrict the integral in (6.13) to the full measure set $X^{\prime}$ where $\|\phi(x)\|\leq\|\phi\|_{L^{\infty}(X)}$ and $|f_{i}(x)|\leq\|f_{i}\|_{L^{\infty}(X)}$ for all $i\in[k]$. Combining Theorem 4.2 and the assumption (6.10) (and assuming as we may that $\phi,f_{1},\ldots,f_{k}$ are supported on $[-CN^{d},CN^{d}]$ for some $C\ll 1$), we see that there is some $\gamma=\gamma(P_{1},\ldots,P_{k})>0$ such that for all $x\in X^{\prime}$ we have the bound
$\left|\frac{1}{(N^{\prime})^{d+1}}\sum_{m\in[(N^{\prime})^{d}],n\in[N^{\prime}]}\theta(n)\phi(T^{m}x)f_{1}(T^{m+P_{1}(n)}x)\cdots f_{k}(T^{m+P_{k}(n)}x)\right|$
$\quad\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N^{\prime})^{-\gamma}$
$\quad\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N)^{-\gamma/2}.$
Now the desired bound (6.13) follows with $c=\gamma/2$. ∎
Finally, we shall prove Theorem 1.3. This follows from the next theorem and Lemma 5.1(ii).
Theorem 6.6 (Pointwise convergence of polynomial ergodic averages with nice weight along a subsequence).
Let $K\in\mathbb{N}$. Let $\theta\colon\mathbb{N}\to\mathbb{C}$ and suppose that there exist some infinite sets $S_{1},\ldots,S_{K}\subset\mathbb{N}$ such that, for all large enough $N\geq 1$ and any $n\in[N]$, we have
(6.14) $|\theta(n)|\leq\sum_{j=1}^{K}\frac{N}{|S_{j}\cap[N]|}1_{S_{j}}(n).$
Suppose also that for all $s\in\mathbb{N}$ and $M\geq 3$ we have
(6.15) $\|\theta\|_{U^{s}[M]}\ll_{s}(\log\log M)^{-\gamma_{s}}$
for some $\gamma_{s}>0$.
Let $N_{j}\gg\exp(\exp(\exp((\log j)^{2}))$ for $j\in\mathbb{N}$. Let $(X,\nu,T)$ be a measure-preserving system. Let $k\in\mathbb{N}$, let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients, and let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$. Then, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$ and $j\geq 1$, we have
$\lim_{j\to\infty}\frac{\exp((\log j)^{3/2})}{N_{j}}\sum_{n\leq N_{j}}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)=0$
for almost all $x\in X$.
This is a direct consequence of the following $L^{1}$ estimate.
Proposition 6.7.
Let the notation and assumptions be as in Theorem 6.6. Then, for all $N\geq 3$, we have
(6.16) $\begin{split}&\|\mathbb{E}_{n\leq N}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)\|_{L^{1}(X)}\\ &\quad\ll_{A,K,P_{1},\ldots,P_{k}}((1+\|f_{1}\|_{L^{q_{1}}(X)})\cdots(1+\|f_{k}\|_{L^{q_{k}}(X)}))^{B(q_{1},\ldots,q_{k})}(\log\log N)^{-c}\end{split}$
for some $c=c(P_{1},\ldots,P_{k})>0$ and $B(q_{1},\ldots,q_{k})>0$.
Proof of Theorem 6.6 assuming Proposition 6.7.
This is immediate from Lemma 5.4(iii). ∎
Proof of Proposition 6.7.
By the same argument as in the proof of Proposition 6.5, we may reduce matters to the case of bounded functions.
Assume now that $q_{1}=\ldots=q_{k}=\infty$. Let $P_{1},\ldots,P_{k}$ be fixed polynomials with integer coefficients with highest degree $d\in\mathbb{N}$. Let $(X,\nu,T)$ be a measure-preserving system, and let $f_{1},\ldots,f_{k}\in L^{\infty}(X)$. We write $A_{N}(x)\coloneqq\mathbb{E}_{n\leq N}\theta(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)$.
By Cauchy–Schwarz, it suffices to show that
$\|A_{N}\|_{L^{2}(X)}\ll\|f_{1}\|_{L^{\infty}(X)}^{2}\cdots\|f_{k}\|_{L^{\infty}(X)}^{2}(\log\log N)^{-c}$
for some $c=c(P_{1},\ldots,P_{k})>0$.
The assumption (6.14) and Lemma 5.5 give
$\|A_{N}\|_{L^{\infty}(X)}\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)},$
so it suffices to show that
$\int_{X}\phi(x)A_{N}(x)\textnormal{ d}\nu(x)\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N)^{-c}$
for all $\phi\in L^{\infty}(X)$ and $N\geq 3$. By the $T$-invariance of $\nu$, this is equivalent to
$\int_{X}\frac{1}{N^{d+1}}\sum_{m\in[N^{d}],n\in[N]}\theta(n)\phi(T^{m}x)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)\textnormal{ d}\nu(x)$
$\quad\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log\log N)^{-c}$
for all $\phi\in L^{\infty}(X)$. But this estimate follows from Theorem 4.2 combined with the assumption (6.15). ∎
Lastly, we prove Theorem 1.5.
Proof of Theorem 1.5.
Following the proof of Theorem 6.4 verbatim, we reduce matters to showing that
(6.17) $\begin{split}&\left|\frac{1}{N^{4}}\sum_{m_{1},m_{2},m_{3},n\in[N]}\mu(n)\phi(T_{1}^{m_{1}}T_{2}^{m_{2}}T_{3}^{m_{3}}x)f_{1}(T_{1}^{m_{1}+n}T_{2}^{m_{2}}T_{3}^{m_{3}}x)\cdots f_{3}(T_{1}^{m_{1}}T_{2}^{m_{2}}T_{3}^{m_{3}+n}x)\right|\\ &\quad\ll_{A}\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{3}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log N)^{-A}\end{split}$
for any functions $\phi,f_{1},f_{2},f_{3}\in L^{\infty}(X)$ and any $x\in X$ for which $|\phi(x)|\leq\|\phi\|_{L^{\infty}(X)},|f_{j}|\leq\|f_{j}\|_{L^{\infty}(X)}$. By setting
$F_{j}((m_{1},m_{2},m_{3}))$ $=f_{j}(T_{1}^{m_{1}}T_{2}^{m_{2}}T_{3}^{m_{3}}x),$
$G((m_{1},m_{2},m_{3}))$ $=\phi(T_{1}^{m_{1}}T_{2}^{m_{2}}T_{3}^{m_{3}}x),$
the estimate (6.17) can be rewritten as
$\left|\frac{1}{N^{4}}\sum_{\mathbf{m}\in[N]^{3},n\in[N]}\mu(n)G(\mathbf{m})F_{1}(\mathbf{m}+(n,0,0))F_{2}(\mathbf{m}+(0,n,0))F_{3}(\mathbf{m}+(0,0,n))\right|$
$\quad\ll_{A}\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{3}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log N)^{-A}.$
But this follows directly from Lemma 4.7. ∎